### v1

# Meta-DAN: towards an efficient prediction strategy for page-level handwritten text recognition

Abstract
Recent advances in text recognition led to a paradigm shift for page-level recognition, from multi-step segmentation-based approaches to end-to-end attention-based ones. However, the naïve character-level autoregressive decoding process results in long prediction times: it requires several seconds to process a single page image on a modern GPU. We propose the Meta Document Attention Network (Meta-DAN) as a novel decoding strategy to reduce the prediction time while enabling a better context modeling. It relies on two main components: windowed queries, to process several transformer queries altogether, enlarging the context modeling with near future; and multi-token predictions, whose goal is to predict several tokens per query instead of only the next one. We evaluate the proposed approach on 10 full-page handwritten datasets and demonstrate state-of-the-art results on average in terms of character error rate. Source code and weights of trained models are available at https://github.com/FactoDeepLearning/meta_dan.
Index Terms:
Handwritten Text Recognition, multi-token prediction, Document Attention Network, full-page.
I Introduction
Handwritten Text Recognition (HTR) is the process of automatically extracting handwritten text from a digitized image and converting it into machine-readable text. The historical approach to tackle this task, still in the majority today, consists of a multi-stage paradigm. A document image is segmented into sections of text, which are then ordered so as to preserve a coherent reading order. Thereafter, these text sections are recognized and reassembled to obtain the overall transcription of the document. Over time, the nature of the textual section has broadened, from the isolated character [1, 2], to the word [3, 4], then to the line [5, 6, 7] and more recently to the paragraph [8, 9, 10, 11, 12], gradually increasing the context usable during recognition. However, regardless of the granularity used, this multi-step approach comes with significant drawbacks: 1) the segmentation step relies on a neural network that needs to be trained, requiring additional segmentation annotations, 2) the ordering step is generally a hand-crafted algorithm that needs to be adapted to the layout of each target dataset, 3) the recognition of the individual sections is carried out independently, limiting the context modeling during prediction, 4) errors accumulate in each of these steps. Also, it has to be noted that works of the literature generally focus on the segmentation or recognition task separately, so that the given recognition results are based on a human-generated segmentation, which is not representative of the task in real-life conditions.
More recently, the Document Attention Network (DAN) [13] was proposed as the first end-to-end approach for HTR at page level, achieving similar recognition performance compared to the multi-step approach. In addition to text, the DAN also predicts special XML-like layout tokens, leading to enriched serialized representations of documents. The DAN architecture includes a transformer-based decoder to perform a character-level autoregressive prediction process. Indeed, the transformer’s attention mechanism is used to generate an implicit segmentation of characters, one after the other, making the model learn the reading order of the document through a text-only supervision. Due to its state-of-the-art performance, the DAN is currently the default approach used for end-to-end handwritten text recognition at page-level [14, 15]. In addition, it has been extended to handle other tasks at page level such as named entity recognition [16, 17, 18], table recognition [19], and key-value information extraction [20, 21, 22], by adding dedicated tokens in the predictable token set.
The main drawback of the DAN, which is also an advantage, lies in its character-level iterative prediction process, which is slow when processing whole pages. Indeed, on the one hand, processing text at character level reduces the vocabulary used, making training more easy. It also makes it possible to manipulate lower-dimensional latent states, since each token carries little information, reducing computational complexity. On the other hand, it increases the size of the generated sequences, and thus, the number of iterations to be performed and the prediction time.
In this work, we propose the Meta Document Attention Network (Meta-DAN), a novel prediction strategy to improve both inference speed and recognition accuracy for end-to-end page-level handwritten text recognition. It lies on the combined use of two distinct approaches. The first one is based on multi-token predictions as it showed interesting results for Natural Language Processing (NLP) [23]. The second consists in windowed queries, i.e., processing successive transformer queries altogether as a sliding window to process them in parallel.
The main contributions of our work are summarized as follows:
  • We propose two prediction strategies to improve page-level handwritten text recognition: windowed queries and multi-token predictions.
  • We study the introduction of a specific dynamic prediction policy to strike a balance between prediction speed and recognition performance for multi-token predictions.
  • We provide an extensive evaluation over a set of 10 datasets, covering several languages and alphabets. It highlights that combining both proposed approaches (Meta-DAN) leads to on-average improved recognition performance compared to concurrent approaches.
  • We show that multilingual page-level text recognition can be beneficial to increase the recognition performance for the mainly represented languages.
II Related Works
II-A Handwritten text recognition
Over the last decades, the field of handwriting recognition has made major advances, moving from isolated character recognition based on expert models, to end-to-end recognition of full-page documents. These advances are mainly due to the modeling capability of deep neural networks, which represent the state of the art in this field. The approaches proposed in the literature can be divided into two main categories we refer to as the alignment-based and the attention-based approaches.
The alignment-based approaches consider the recognition task as a one-dimensional sequence alignment problem between a prediction sequence and a ground truth sequence. Modeling the text recognition task in this way enables the end-to-end recognition of word [24, 25] and line [5, 7] images. The practical application of this method is based on two aspects: the reduction of the vertical dimension, generally achieved through convolution or pooling, and the use of the Connectionist Temporal Classification (CTC) [26] introduced in 2006. The CTC loss enables to align a probability lattice computed from the input image with a sequence of characters corresponding to the ground truth. The strength of this approach is also its weakness: it only deals with one-dimensional sequences. To go a step further, two works showed that it is feasible to recognize whole paragraphs by using tricks to assimilate a paragraph to a long line of text, through bilinear interpolations [10], or concatenation [27]. However, to the best of our knowledge, this alignment-based approach is not suited for end-to-end full-page recognition, since the complexity of page layouts (e.g., multi-column pages) makes it impossible to reduce it to a one-dimensional sequence alignment problem.
More recently, the attention-based approaches emerged, particularly with the rise in popularity of the Transformer architecture [28]. They rely on attention mechanisms to provide an iterative decoding process at character level, performing a kind of implicit character segmentation through the attention maps. In other words, characters are predicted one after the other following an autoregressive process. At each decoding step, attention weights are computed and matched to the extracted features of the input image, with the aim of focusing on a specific part of the image (the character to recognize). Contrary to the alignment-based approaches, the attention-based approaches can be applied to any input image, from a single line [6, 29] or paragraph [9, 30, 11] to a two-page document with the DAN-based works [13, 14, 17]. The main drawback of such approach lies in the prediction time that grows with the length of the text sequence to be recognized.
It has to be noted that marginal works [8, 12] were proposed to handle paragraph recognition with a hybrid approach combining the attention mechanisms, to generate an implicit line segmentation, and the alignment-based approach, to recognize them iteratively through CTC decoding. In addition, an attempt of adapting the CTC to 2D-to-1D alignment was proposed in [31], but it did not lead to convincing results.
In summary, alignment-based approaches are advantageous because of their one-shot prediction, but they are limited to images with a single column of text, whereas attention-based approaches are capable of handling entire pages with complex layout but involve longer prediction times. Attention-based decoding strategies also suffer from a limited context modeling, in that it only has access to the previously predicted tokens to make its prediction (only past is considered). In this work, we focus on reducing the prediction time of the attention-based character-level approach to reach fast end-to-end page-level predictions, while improving the context modeling through the integration of near future in the decoding process.
II-B Speeding up autoregressive processes
The prediction time issue of autoregressive processes can be found in various applications including a text generation stage, and several strategies have been proposed. In [32], a transformer-based network is trained to predict the $n$ future tokens instead of a single one in the context of text summarization. We refer to this strategy as multi-token prediction. This idea was successfully applied for large language models recently [23]. In this latter work, $k$ prediction heads are applied on the same transformer output, each one predicting one of the $k$ next tokens, limiting the computational overhead. A similar approach was proposed for neural machine translation in [33] by including an extra verification step: prediction is performed another time using all the predictions as in teacher forcing to validate them or not.
The authors of [34] proposed another way to reduce the prediction time for the task of image captioning. The target sequence is split into subgroups of words, which are decoded in parallel. They use an extra DEL token to remove duplicated predictions. Closer to our work, a similar approach was proposed to speed up the Document Attention Network by splitting the decoding process into two stages [14]. The first stage is dedicated to predicting the first character of each line, and the second one consists in completing all lines in parallel, reducing the number of decoding steps. While the prediction time is significantly decreased, this alternative suffers from a quite chaotic context: the model has access to the beginning of each line, i.e., a partial view of both past and future. The windowed queries we propose also rely on the idea of processing queries in parallel, but they differ in that the processed queries are successive, leading to a unique decoding stage and a context mixing full past and near future.
An alternative to speed up the prediction process is to reduce the length of the target sequence. This can be achieved through a specific tokenizer, one of the best known is the Byte-Pair Encoding (BPE) algorithm [35]. Instead of encoding the text to be recognized as a sequence of character tokens, it is encoded as tokens representing sub-words of variable sizes, obtained from their frequency of occurrence in a given dataset. This technique has been applied in [17] showing a significant prediction time decrease, in return for a higher character error rate. The approach we propose is compatible with this kind of technique as it rely on the token prediction strategy, regardless of the nature of the tokens.
In this paper, we propose to combine both windowed queries and multi-token predictions to speed up the prediction process while improving the context modeling capability of the model. While evaluated for the HTR task in this paper, the approach is generic and could be applied for any task relying on a transformer-based iterative decoding process.
III Background
DAN architecture. Encoding stage: 2D features $\boldsymbol{f}$ are extracted from the input image through a Fully Convolutional Network (FCN). Features are enhanced with 2D positional encoding and flattened for transformer needs. Decoding stage: at each iteration $t$, a query sequence $\boldsymbol{q}^{t}$ is generated from the previous predictions $\hat{\boldsymbol{y}}^{t}$ and associated to the image features through transformer’s attention mechanisms to compute the new prediction $\hat{\boldsymbol{y}}_{t}$.
To make things clearer, we first propose a formulation of the handwritten text recognition problem, we go over the main points of the DAN [13] approach on which is built this work, and we discuss the Faster DAN [14] proposition.
III-A Problem formulation
[…]
III-B Document Attention Network
The Document Attention Network (DAN) [13] was proposed to tackle this task as a character-level autoregressive prediction process. While the original approach generates additional layout tokens in addition to text tokens, we only focus on text recognition in this work. The DAN refers to a deep neural network illustrated in Figure III. It is made up of two main building blocks: a Fully Convolutional Network (FCN) encoder and a causal transformed-based decoder. The general idea of the approach is as follows. Based on the image features extracted by the encoder, and on the previously predicted tokens, the goal is to predict the next token (each token representing a character).
The DAN encoder, noted $e$, was first proposed in [12] as a generic feature extractor for handwritten text recognition able to deal with images of variable sizes. It mainly consists in a succession of 18 standard convolutions and 12 Depthwise Separable Convolutions [36], coupled with instance normalization, mix dropout layers and skip connections. It extracts 2D features $\boldsymbol{f}=e(\boldsymbol{X})$ ($\boldsymbol{f}\in\mathbb{R}^{H_{f}\times W_{f}\times C_{f}}$) from the input image by reducing the height by 32 and the width by 8 ($H_{f}=\frac{H}{32}$, $W_{f}=\frac{W}{8}$ and $C_{f}=256)$. All details can be found in the original paper [12].
[…]
For each iteration $t$, the previously predicted tokens are put together as query input for the decoder: $\hat{\boldsymbol{y}}^{t}=[\hat{\boldsymbol{y}}_{0},\dots,\hat{\boldsymbol{y}}_{t-1}]$. A trainable weight matrix is used to embed each text token as a vector of dimension $C_{f}$, leading to $\boldsymbol{e}\in\mathbb{R}^{t\times C_{f}}$. Following the original transformer [28], one-dimensional positional encoding ($\boldsymbol{p}^{t}_{\text{1D}}$) is added, leading to the query sequence $\boldsymbol{q}^{t}=\boldsymbol{e}^{t}+\boldsymbol{p}^{t}_{\text{1D}}$. In the same way, its 2D adaptation [30] encoding both dimensions of the image on half of the channels is used to inject positional information to the image features, resulting in $\boldsymbol{g}=\boldsymbol{f}+\boldsymbol{p}_{\text{2D}}$. Query tokens $\boldsymbol{q}^{t}$ and images features $\boldsymbol{g}$ are processed through 8 stacked transformer decoder layers resulting in $\boldsymbol{o}^{t}=d(\boldsymbol{q}^{t},\boldsymbol{g})$, preserving the number of channels $C_{f}$. Finally, a score $\boldsymbol{s}_{t}$ for each token of the token set $\mathcal{A}$ is computed through a linear projection layer $p$ based on the output corresponding to the last token of the query sequence:
$\boldsymbol{s}_{t}=p(\boldsymbol{o}_{t-1}).$ (1)
The predicted token corresponds to the one with the highest score: $\hat{\boldsymbol{y}}_{t}=\operatorname*{arg\,max}(\boldsymbol{s}_{t})$.
The DAN is trained using the cross-entropy loss $\mathcal{L}_{\text{CE}}$ over the sequence of logits of the scores obtained through softmax activation:
$\mathcal{L}_{\text{DAN}}=\sum_{t=1}^{L+1}\mathcal{L}_{\text{CE}}(\boldsymbol{y}_{t},\boldsymbol{p}_{t}),$ (2)
with $\boldsymbol{p}_{t,i}=\operatorname*{soft\,max}(\boldsymbol{s}_{t})=\frac{e^{\boldsymbol{s}_{t,i}}}{\sum_{j}e^{\boldsymbol{s}_{t,j}}}$.
[…]
[width=]images/tikz/dan_in_out DAN [13]
[…]
[width=]images/tikz/fdan_in_out Faster DAN [14]
[…]
[width=]images/tikz/wdan_in_out W-DAN (ours, $w=3$)
[…]
[width=]images/tikz/mt-dan_in_out MT-DAN (ours, $m=3$)
[…]
[width=0.928]images/tikz/meta_dan_in_out Meta-DAN (ours, $w=3$, $m=2$)
[…]
III-C Faster DAN
The Faster DAN [13] was proposed as an alternative to speed up the decoding process at inference stage. It has to be noted that the Faster DAN architecture is exactly the same as the DAN. The contribution is about decoding strategy whose aim is to parallelize the text line predictions. It relies on a two-stage approach. First, an autoregressive process predicts the first character of each line, following the original DAN strategy. Then, all lines are completed in parallel, reducing the total number of iterations. The main difference appears in that second decoding stage. Instead of using the output corresponding to the last token $\boldsymbol{o}_{t-1}$ only, the token prediction is performed using the outputs corresponding to the last token of each text line. In this way, for $N$ text lines, there are $N$ token predictions per iteration. Equation 1 becomes:
$\boldsymbol{s}^{l}_{t}=p(\boldsymbol{o}^{l}_{t-1}),$ (3)
where $\boldsymbol{s}^{l}_{t}$ is the score vector associated to the token prediction of the next character for line $l$ at iteration $t$, and $\boldsymbol{o}^{l}_{t-1}$ is the transformer output corresponding to the last predicted token for line $l$.
The Faster DAN introduces a specific positional encoding to associate each text token with a line index and a positioning index within that line. This is a crucial component to make this parallelization possible. Although the Faster DAN showed interesting inference speed improvements, it results in a loss of past context when predicting a character. Indeed, instead of benefiting from the whole past context as is the case for the DAN, the Faster DAN has a partial view of the past and future due to the parallel line predictions, which reduces its ability to model the language.
In this paper, we propose a strategy to get the best of both worlds: a reduced prediction time through query parallelization, and an improved context modeling by processing successive queries altogether and relying on multi-token predictions.
IV Meta Document Attention Network
We propose two new approaches for the autoregressive prediction of text tokens in the context of end-to-end handwritten text recognition at page level. For both, the idea is to predict multiple tokens at each decoding iteration to reduce the number of decoding steps, while preserving the whole past context. We assume this is a key element to reduce the number of recognition errors through language modeling. First, we propose to associate multiple predictions to a single query $q_{t-1}$, to predict the $m$ next tokens at once. This is referred as the Document Attention Network with Multi-Tokens predictions (MT-DAN). Second, we propose a query approach based on a window mechanism to process multiple queries at once, each associated with a single prediction. This is referred as the Document Attention Network with Windowed queries (W-DAN). Finally, we design a global strategy including these two approaches, called Meta Document Attention Network (Meta-DAN).
A visual comparison of the prediction processes is depicted in Figure III-B.
IV-A MT-DAN
Inspired by recent works on multi-token predictions for Natural Language Processing (NLP) [23], we propose the Document Attention Network with Multi-Tokens predictions (MT-DAN). It relies on the following process. Based on the last query output $\boldsymbol{o}_{t-1}$, $m$ projection heads are used to compute the scores for the $m$ next tokens, with $m$ being a new hyperparameter. Equation 1 becomes:
$\boldsymbol{s}^{(\boldsymbol{t})}_{t+k}=p^{k}(\boldsymbol{o}_{t-1}),$ (4)
where $\boldsymbol{s}^{(\boldsymbol{t})}_{t+k}$ is the score vector associated to the token prediction of the $k^{\text{th}}$ next token computed at iteration $t$, $p^{k}$ is the $k^{\text{th}}$ projection head, and $\boldsymbol{o}_{t-1}$ is the transformer output corresponding to the last query token $\boldsymbol{q}_{t-1}$. In this way, $m$ predictions are performed per iteration following:
$\hat{\boldsymbol{y}}^{(\boldsymbol{t})}_{t+k}=\operatorname*{arg\,max}(\boldsymbol{s}^{(\boldsymbol{t})}_{t+k}),$ (5)
with $k\in[0,\dots,m-1]$.
As one can note, several predictions $\hat{\boldsymbol{y}}^{(\boldsymbol{t})}_{t+k}$ will be associated to the same target token $\boldsymbol{y}_{t+k}$. This is not an issue during training, we train the MT-DAN using the cross-entropy loss over the whole token sequence, as for the original DAN. It involves adding $m-1$ extra <e> tokens in order for all text tokens to have a target for each of the projection heads, leading to the following loss:
$\mathcal{L}_{\text{MT-DAN{}}}=\sum_{t=1}^{L+1}\sum_{k=0}^{m-1}\mathcal{L}_{\text{CE}}(\boldsymbol{y}_{t+k},\boldsymbol{p}^{(\boldsymbol{t})}_{t+k}),$ (6)
with $\boldsymbol{p}^{(\boldsymbol{t})}_{t+k}=\operatorname*{soft\,max}(\boldsymbol{s}^{(\boldsymbol{t})}_{t+k})$.
In Section VI, we propose a decoding policy based on the confidence level of the prediction to choose not to predict all $m$ next tokens for poor confidence. We demonstrate it is a good trade-off between prediction speed and recognition performance.
Predicting multiple tokens $\hat{\boldsymbol{y}}_{t+k}$ from a unique output $\boldsymbol{o}_{t-1}$ can theoretically lead to saturation as $m$ increases. Indeed, the fixed-size vector $\boldsymbol{o}_{t-1}$ must represent the next $m$ text tokens at once. To solve this issue, we propose another way to predict multiple tokens per iteration by processing several consecutive queries at once.
IV-B W-DAN
[…]
[width=]images/tikz/dan_attention DAN [13]/MT-DAN (ours)
[…]
[width=]images/tikz/fdan_attention Faster DAN [14]
[…]
[width=]images/tikz/wdan_attention W-DAN (ours) / Meta-DAN (ours)
(for $w=3$)
[…]
To solve the MT-DAN output compression issue, we propose the Document Attention Network with Windowed queries (W-DAN). Instead of relying on a single output to predict $m$ tokens, we opt for using the output of the last $w$ predictions to predict the $w$ next tokens, $w$ being a new hyperparameter. In this way, it solves the compression issue as they are as many output as they are predictions. More precisely, each output $\boldsymbol{o}_{t-q}$ is responsible for the prediction of token $\hat{\boldsymbol{y}}_{t-q+w}$, with $q\in[1,\dots,w]$. Equation 1 becomes:
$\boldsymbol{s}_{t-q+w}=p(\boldsymbol{o}_{t-q}),$ (7)
with
$\hat{\boldsymbol{y}}_{t-q+w}=\operatorname*{arg\,max}(\boldsymbol{s}_{t-q+w}).$ (8)
[…]
$\mathcal{L}_{\text{W-DAN{}}}=\sum_{t=w}^{L+1+n_{\text{e}}}\mathcal{L}_{\text{CE}}(\boldsymbol{y}_{t},\boldsymbol{p}_{t}),$ (9)
There is a major advantage of both presented approaches compared to the DAN and the Faster DAN: the $k$ or $w$ predictions follow one another and are performed altogether. It means that the context is enlarged when prediction takes place. The model must focus on a broader context to predict multiple successive tokens. For instance, for $k=3$ (or $w=3$), the model focuses on predicting $\{\hat{\boldsymbol{y}}_{t},\hat{\boldsymbol{y}}_{t+1},\hat{\boldsymbol{y}}_{t+2}\}$ at once. This way, the prediction of $\hat{\boldsymbol{y}}_{t}$ is impacted by the future context $\{\hat{\boldsymbol{y}}_{t+1},\hat{\boldsymbol{y}}_{t+2}\}$, which improves the ability to model the language at prediction time compared with DAN (which only benefits from past context). The same goes with $\hat{\boldsymbol{y}}_{t+1}$ predicted with $\{\hat{\boldsymbol{y}}_{t+2}\}$ as context of future.
IV-C Meta-DAN
We propose now to formulate a generic prediction strategy to take advantage of both presented approaches by combining them. It is referred as the Meta Document Attention Network (Meta-DAN). Thus, the Meta-DAN is parametrized by two hyperparameters $m$ and $w$, indicating the number of projection heads and the window size, respectively. In other words, at each iteration $t$, a window of $w$ query tokens $\{\boldsymbol{q}_{t-w},\dots,\boldsymbol{q}_{t-1}\}$ are processed at once, each one leading to $m$ predictions. More precisely, the output of the query token $\boldsymbol{q}_{t-q}$, i.e. $\boldsymbol{o}_{t-q}$, is responsible for the prediction of $\{\hat{\boldsymbol{y}}_{t-q+w},\dots,\hat{\boldsymbol{y}}_{t-q+w+m-1}\}$ as follows:
$\hat{\boldsymbol{y}}^{(\boldsymbol{t})}_{t}=\operatorname*{arg\,max}(\boldsymbol{s}^{(\boldsymbol{t})}_{t}),$ (10)
with
$\boldsymbol{s}^{(\boldsymbol{t-q+1})}_{t-q+w+k}=p^{k}(\boldsymbol{o}_{t-q}),$ (11)
where $\boldsymbol{s}^{t-q+1}_{t-q+w+k}$ is the score associated to token $\hat{\boldsymbol{y}}_{t-q+w+k}$ by the $k^{\text{th}}$ projection head of the $q^{\text{th}}$ query token in the query window, and with $k\in[0,\dots,m-1]$ and $q\in[1,\dots,w]$. The Meta-DAN involves initializing the decoding process with $w$ <s> tokens, as for the W-DAN, and adding extra <e> so that the last query has a target prediction for all its associated projection heads.
Regarding training, equation 2 becomes:
$\mathcal{L}_{\text{Meta-DAN{}}}=\sum_{t=w}^{L+1+n_{\text{e}}}\sum_{k=0}^{m-1}\mathcal{L}_{\text{CE}}(\boldsymbol{y}_{t+k},\boldsymbol{p^{(\boldsymbol{t})}_{t+k}})$ (12)
Detailed features of evaluated datasets (page-level). Dataset Train Val Test Charset Average Language Century Writers size # char. per page IAM (paragraph) [37] 747 161 336 80 371 English 20th-21st 657 RIMES-2009 [38] 1,050 100 100 108 588 French 21st $\approx 400$ RIMES-2024 [39] 4,500 505 600 123 567 French 21st $\approx$ 1,400 READ-2016 [40] 350 50 50 89 468 German 15th-19th unknown Esposalles [41] 75 25 25 62 1,748 Catalan 17th 1 BRESSAY [42] 647 154 199 109 2,612 Portuguese 21st 1,000 ScribbleLens [43] 164 14 21 85 1,410 Dutch 16th-18th 85 Eparchos [44] 80 20 20 157 962 Greek 16th 2 CASIA-2 [45] 3,600 475 1,015 2,704 273 Chinese 21st 1,019 MAURDOR-C3 [46] 1,006 148 166 134 491 French/English 21st > 130 MAURDOR-C4 [46] 721 111 114 127 675 French/English 21st > 90
As one can note, several query tokens can be in charge of the same prediction. This is shown in Figure III-B with a Meta-DAN example with $w=3$ and $m=2$: the "m" and "p" are predicted twice, from two different queries. In this work, we limit ourselves to use the prediction of the first projection head for the $w-1$ first queries, and the prediction of all projection heads for the last query. While discarded, we assume that training to predict those duplicated characters encourages the network to use its future context to improve language modeling. In this way, all queries follow the same behavior, which make training easier.
In summary, the Meta-DAN approach is a general formulation for which one can distinguish three specific cases depending on the hyperparameters:
  • Choosing $m=1$ and $w=1$ consists in the DAN approach.
  • Choosing $m>1$ and $w=1$ results in the MT-DAN approach.
  • Choosing $m=1$ and $w>1$ leads to the W-DAN approach.
IV-D Impact on the computational complexity
To further give insights about the main differences between all these prediction strategies, Figure IV-B provide visuals of the attending maps (transformer masks), stating which token can access to which other tokens.
Regarding the MT-DAN approach, the attending maps of the DAN and the MT-DAN are identical; this means that they both imply the same computations in the transformer decoder layers. In addition, contrary to the NLP application which relies on the Byte-Pair Encoding algorithm leading to large vocabularies, the addition of extra projection heads do not significantly impact the global computational complexity as we are working at character level, limiting the dimensions of the projection weight matrices.
Concerning the W-DAN approach (and thus the Meta-DAN), the attending maps show that there are more interactions between the tokens in the self-attention layers, resulting in more computations, theoretically. However, to our knowledge, the current implementations of the transformer relies on masks to "hide" the forbidden tokens (usually the future) after the query-key multiplication takes place. In this way, this does not impact the computation time in practice since computations are performed over all token pairs whether they are really used or not.
No matter the approach (DAN, Faster DAN, MT-DAN, W-DAN, Meta-DAN), the prediction process can be parallelized at training time using teacher forcing, so there is no additional benefit for the training stage. As stated previously, the number of computations are not significantly impacted neither by the MT-DAN nor by the W-DAN. The major impact is positive and related to the time complexity at prediction time when parallelizing the computation of token predictions.
V Experimental conditions
V-A Datasets
We evaluate the proposed approach on many public HTR datasets, namely IAM [37], READ-2016 [40], Esposalles [41], BRESSAY [42], ScribbleLens [43], Esparchos [44], CASIA-2 [45], MAURDOR [46], RIMES-2009 [38] and RIMES-2024 [39]. They cover 8 languages, three alphabets (Latin, Greek and Chinese), 6 centuries (15th-21st). It enables us to have a robust evaluation due to the high diversity of the datasets. Indeed, as detailed in Table IV-C, they involve many writing configurations (from single writer to more than a thousand), various sizes of character set, as well as a variable amount of training data (from less than a hundred to few thousand pages).
We used the standard split into training, validation and test sets when existing. To our knowledge, we are the first to use the recently released RIMES-2024 dataset, so we propose a new split, randomly selecting samples for training, validation and test by respecting the split provided at line level in 2011 [47]. We discarded images from the "Q" category, which corresponds to forms, since they do not provide the whole transcript. For multilingual experiments, we discard the RIMES-2009 dataset from the evaluation setting as some test samples can be found in the RIMES-2024 training set. For ScribbleLens, we removed 10% of the training samples to create a validation set by randomly picking from each of its data sources. For Eparchos, we propose a split at page-level by taking 80% of the samples for training, 10% for validation and 10% for test. Appendix B provides the splits at line level, used for pre-training purposes, for the datasets providing line-level annotations. This section also includes some input document image examples.
[b] Comparison of W-DAN recognition performance with respect to the window size $w$ on the IAM and BRESSAY test sets. Speed Factor (SF) is indicated as a comparison with the DAN baseline.
Window IAM BRESSAY
size ($w$) CER (%) WER (%) Time (s) SF CER (%) WER (%) Time (s) SF
1 (DAN) 4.23 14.00 3.39 x1.0 2.43 7.85 24.65 x1.0
2 4.67 15.14 1.94 x1.7 2.91 8.87 20.86 x1.2
3 4.83 15.62 1.26 x2.7 3.05 9.24 12.25 x2.0
4 4.75 15.71 0.97 x3.5 2.63 9.35 8.63 x2.8
5 4.87 16.28 0.77 x4.4 3.02 9.72 6.63 x3.7
10* 5.07 16.53 0.40 x8.5 2.66 9.53 2.95 x8.4
25* 6.56 20.94 0.18 x18.8 4.37 15.21 1.13 x21.8
    Training was performed over 2M and 3M training samples for $w=10$ and $w=25$, respectively (instead of 1M).
V-B Training strategy
If not stated otherwise, we follow the main guidelines described in [13]:
  • We pre-train the encoder part of the architecture using a line-level training strategy. An adaptive max-pooling layer is added on top of the encoder to collapse the vertical axis, leading to a one-dimensional sequence of features. Then classification layer is applied to generate a probability lattice. This pre-training is carried out using the CTC loss. If line-level annotations are provided in the dataset, we use a ratio of 25% real / 75% synthetic images (100% synthetic otherwise).
  • We randomly apply some data augmentation strategies with a probability of 90%. It includes resolution modification, color jittering, Gaussian noising, Gaussian blurring, elastic distortion, dilation, erosion and perspective transformation.
  • During page-level training, we rely on a curriculum synthetic document generation strategy. The generated documents are more and more complex through the training, starting with single line images to documents including several dozens lines. The model is trained on both real and synthetic documents with the proportion of synthetic documents slowly decreasing from 90% to 10%. Following [17], we scraped the website 1001-fonts.com to get a large collection of fonts from which we randomly sample when rendering a synthetic text line. The text used for synthetic generation comes from the training set of the target dataset itself.
  • Training and pre-training are performed using the Adam optimizer with an initial learning rate of $10^{-4}$, and automatic mixed-precision.
  • Line-level pre-training is carried out using 5M training samples, and page-level training (including curriculum) is performed over 1M samples.
  • Input images are preprocessed to reach a resolution of 150 dpi by preserving the original aspect ratio, whether it is for pre-training, training or evaluation.
  • We do not use any external data, external language model nor lexicon constraints.
V-C Evaluation
We use the standard Character Error Rate (CER) and the Word Error Rate (WER) to evaluate the text recognition performance. Both are computed with the Levenshtein distance between the ground truth and the predicted text, normalized by the length of the ground truth. It has to be noted that page-level approaches are sensitive to reading order; it means that correct text recognized in a wrong order will be considered as incorrect. For each model, we compute the average prediction time for one test image when using a single A100 GPU and a mini-batch size of 1. To provide fully comparable experiments, notably between monolingual and multilingual experiments we present in the following, we set the same maximum number of decoding iterations for all datasets. It is set to 5,000 to cover the requirements of all evaluated dataset. It has to be noted that this choice results in recognition metrics that can highly be impacted by the prediction loops that can occur, i.e., when the model predicts again and again the same text line.
VI Experiments
In this section, we evaluate the proposed approaches (W-DAN, MT-DAN and Meta-DAN) on the IAM and BRESSAY datasets to select the best approach and hyperparameters. While relying on the original DAN architecture [13], we propose an enlarged version of the encoder we refer to as FCN-M as it shows better recognition performance (details in Appendix A). Then, we apply them on the other 9 datasets to validate the robustness of the approach, and compare it to the state of the art. Finally, we present the results for a multilingual recognition setting.
VI-A W-DAN
We evaluate the W-DAN approach for window sizes $w\in\{1,2,3,4,5,10,25\}$, with $w=1$ corresponds to the DAN setting. Results are shown in Table V-A. It highlights that it is possible to predict from 2 to 10 character tokens at once, while still being competitive in terms of CER and WER. It is especially true for the BRESSAY dataset for which $w=10$ leads to a small increase of only 0.23 point of CER, for an 8.4 speed factor. However, results degrade significantly for $w=25$ with a CER of 6.56% instead of 4.23% for IAM, and 4.37% instead of 2.43% for BRESSAY, even with three times more training samples. As one can note, it is possible to reduce the average prediction time per sample nearly linearly with respect to the window size. This is illustrated in Figure VI-A. Regarding the BRESSAY prediction time, it appears that the new implementation has impacted the prediction time for large text predictions, leading to a lag in the speed factor improvement. A main observation is that the model is harder and harder to train as $w$ increases, so we needed to increase the amount of training data for $w=10$ and $w=25$ up to 2M and 3M, respectively.
Evolution of the prediction time with respect to the window size $w$ for the W-DAN and to the number of token heads $m$ for the MT-DAN on the test set of IAM and BRESSAY datasets. The prediction time dramatically decreases as $m$ or $w$ increases at first, and then less sharply as the computation load of the encoder becomes more and more significant on the overall inference process.
VI-B MT-DAN
[b] Comparison of MT-DAN recognition performance with respect to the number of token heads $m$ on the IAM and BRESSAY test sets. Speed Factor (SF) is indicated as a comparison with the DAN baseline. Number of IAM BRESSAY heads ($m$) CER (%) WER (%) Time (s) SF CER (%) WER (%) Time (s) SF 1 (DAN) 4.23 14.00 3.39 x1.0 2.43 7.85 24.65 x1.0 2 4.70 14.85 1.84 x1.8 2.47 8.54 13.02 x1.9 3 4.74 15.17 1.23 x2.8 2.46 8.60 8.85 x2.8 4 4.68 15.56 0.93 x3.6 2.70 8.91 6.64 x3.7 5 5.14 16.01 0.75 x4.5 2.45 8.72 5.28 x4.7 10* 5.06 16.29 0.39 x8.7 2.37 8.64 2.64 x9.3 25* 7.19 23.83 0.16 x21.2 3.15 11.28 1.11 x22.2
    Training was performed over 2M training samples for $m=10$ and $m=25$. An additional character-level curriculum stage was also added for the early stages of the training.
We evaluate the MT-DAN approach in a similar setting, i.e., for various numbers of token heads $m\in\{1,2,3,4,5,10,25\}$, with $m=1$ corresponds to the original DAN. The obtained results (Table VI-B) are quite similar to those of the W-DAN, with comparable recognition performance and prediction time. Neither approach seems to stand out from the other. Even for the extreme case of $m=25$, it appears that the MT-DAN performs better for the BRESSAY dataset with a CER of 3.15% compared to 4.37% for the W-DAN (with $w=25$), but we observe the contrary for the IAM dataset with a CER of 7.19% for the MT-DAN compared to 6.56% for the W-DAN. As for the W-DAN, it required more training data for $m=10$ and $m=25$. We also relied on a curriculum strategy at character level before the line-level curriculum strategy takes place. We did not use it for the W-DAN as it was not effective; we assume this is due to the way the tokens interact for make its predictions in this approach.
Improving the prediction policy
To enhance the capacity of the MT-DAN approach, we propose and evaluate two prediction policies at inference time:
1) The static prediction strategy consists in always predicting the $k$ next tokens at each decoding iteration ($k$ must be lower or equal to the number of tokens heads $m$). Previous results (shown in Table VI-B) relied on this naïve prediction policy, setting $k=m$.
2) The dynamic prediction strategy involves predicting between the next 1 and $m$ tokens, based on a confidence threshold $\tau$. This way, the number of predicted tokens must vary from one decoding iteration to another. The idea is to accept to slow down the decoding process when the model is unsure of its prediction, by assuming that it could be better at the next decoding iteration for which it will benefit from more context. The first token head is always used to avoid infinite loop; the other $m-1$ heads are then considered sequentially. If the confidence score (i.e., the maximum value after softmax operation) associated to the current head is greater or equal to $\tau$, then the prediction is preserved, and the next head is processed; the decoding iteration ends otherwise.
We first study the choice of the confidence threshold $\tau$ for the dynamic strategy. Figure VI shows the evolution of the CER with respect to $\tau$ for values between 0% and 100%, with a step of 5%. We compare it for different numbers of heads and for both IAM and BRESSAY datasets. It appears that the recognition performance can globally be improved as of $\tau=0.5$, except for the configurations $m=3$ and $m=4$ on the BRESSAY dataset whose behavior is less monotonous. One should note that this prediction policy can even reach better CER performance than the DAN (blue line) on the BRESSAY dataset, for $w=2$ and $w=5$. We consider $\tau=0.9$ as a good compromise over all the evaluated configurations and use it for the next experiments.
Comparison of static and dynamic prediction policies of the MT-DAN. CER (%) are given for the test sets of IAM and BRESSAY datasets. Best results are shown in bold; results that are very close to the best (whose difference in values is less than 0.05%) are underlined. Number IAM BRESSAY of heads static dynamic static dynamic ($m$) $k=1$ $k=2$ $k=3$ $k=4$ $k=5$ $\tau=0.9$ time (s) $k=1$ $k=2$ $k=3$ $k=4$ $k=5$ $\tau=0.9$ time (s) 1 (DAN) 4.23 2.43 2 4.33 4.70 4.59 1.92 2.41 2.47 2.36 13.60 3 4.31 4.85 4.74 4.31 1.32 2.46 2.46 2.46 2.69 9.21 4 4.48 4.54 4.68 4.68 4.52 1.00 2.71 2.86 2.61 2.70 2.71 6.97 5 4.45 4.74 4.54 4.86 5.14 4.51 0.83 2.34 2.40 2.36 2.40 2.45 2.33 5.70 10 4.79 4.80 4.87 4.87 4.90 4.80 0.47 2.67 2.24 2.70 2.28 2.29 2.66 3.15 25 5.73 5.61 5.58 5.64 5.68 5.47 0.36 3.35 3.30 2.92 2.91 2.93 2.87 1.88
Impact of the CER threshold for dynamic decoding of the MT-DAN on the test set of IAM and BRESSAY.
We compare the two prediction policies, namely static and dynamic, in Table VI. For the static prediction strategy, we evaluate the models for $k\in\{1,2,3,4,5\}$ when possible. Regarding the IAM dataset, one can observe a trend whereby the character error rate decreases as $k$ also decreases, with optimum performance reached for $k=1$, noting that decreasing $k$ highly increases the prediction time as shown previously. This trend is less noticeable on the BRESSAY dataset with more fluctuating results. However, the proposed dynamic prediction policy with $\tau=0.9$ shows interesting results which generally outperforms the corresponding model using the static strategy with $k=m$, with a limited prediction time overhead. For instance, it enables to reduce the CER from 5.14% to 4.51% for the IAM dataset ($w=5$), while slightly increasing the prediction time from 0.75 second (Table VI-B) to 0.83 second.
VI-C Meta-DAN
In this section, we focus on $m=5$ and $w=5$ as it seems to be a good comprise between prediction time and recognition performance.
Pre-training strategy
Comparison of Meta-DAN ($m=5$, $w=5$) pre-training strategy on the test sets of IAM and BRESSAY. Pre-training curriculum IAM BRESSAY line-level (baseline) 5.59 3.01 W-DAN ($w=5$) 4.82 2.51 W-DAN ($w=5$) 5.01 2.42 MT-DAN ($m=5$) 5.04 2.36 MT-DAN ($m=5$) 4.79 2.38
In Table VI, we evaluate the Meta-DAN for several pre-training strategies. We compare the standard line-level pre-training to transfer learning directly from page-level W-DAN or MT-DAN, with and without curriculum learning. We observe that best results are obtained through page-level pre-training, highlighting the usefulness of the pre-trained decoder even if it does not rely on the same decoding strategy. We also note that re-training from line-level through curriculum learning is generally not necessary to reach the best results. Lastly, it seems that the MT-DAN pre-training leads to slightly better CER on average. For these reasons, the Meta-DAN training will be based on MT-DAN pre-training without curriculum learning in the following experiments.
Comparison with MT-DAN and W-DAN
We compare the Meta-DAN with the MT-DAN and the W-DAN for the same hyperparameters ($m=5$ and $w=5$) in Table VI. As one can note for the IAM and BRESSAY datasets, combining both windowed queries and multi-token predictions enable to increase the recognition performance, as well as reducing the prediction time. Comparing to DAN, the conclusion is not as clear-cut regarding the character error rate as it is in favor to the DAN for IAM with a CER of 4.23% compared to 4.79%, but in favor of the Meta-DAN for BRESSAY with a CER of 2.38% compared to 2.43%. It shows that the Meta-DAN can be both more efficient and faster than the DAN depending on the target dataset.
Comparison of proposed approaches using the static prediction policy (with $m=k$). CER are expressed in percentages, and times in seconds. Approach $m$ $w$ IAM BRESSAY CER Time CER Time DAN 1 1 4.23 3.39 2.43 24.65 W-DAN 1 5 4.87 0.77 3.02 6.63 MT-DAN 5 1 5.14 0.75 2.45 5.28 Meta-DAN 5 5 4.79 0.43 2.38 4.01
Comparing prediction strategies
As the Meta-DAN also relies on multi-token predictions, we can apply the same dynamic prediction policy as for the MT-DAN. Its usage is evaluated in Table VI. It appears that the confidence scores associated to the predictions are not as high as for the MT-DAN, leading to a failure case for the BRESSAY dataset for which the prediction time is close to that of W-DAN (static case with $k=1$). As noted previously, the static policy with $k=1$ leads to the best recognition performance but does not benefit from accelerated prediction. We consider the static policy with $k=m=5$ as the best alternative for the Meta-DAN and use this configuration in the following.
Comparison of Meta-DAN ($w=5$, $m=5$) performance with respect to the decoding strategy on the test sets of IAM and BRESSAY. Strategy for IAM BRESSAY concurrent CER (%) Time (s) CER (%) Time (s) static ($k=1$) 4.55 0.79 2.31 7.16 static ($k=5$) 4.79 0.43 2.38 4.01 dynamic ($\tau=0.9$) 4.73 0.45 2.37 7.14
VI-D Visualization
A visualization of the attention process for the different prediction approaches is provided in Figure VI-D. Displayed images are cropped to a single line for visibility, but predictions are performed on the full-page images. Attention weights of the last cross-attention layer are projected on the input image and values are represented by color intensity (maximum intensity for 0.2 value). One color is associated to each projection head in this order: red, blue, green, yellow and cyan. As one can notice, the attention weights are more peaky for DAN and W-DAN. It was expected as each query only focuses on a single character prediction, contrary to MT-DAN and Meta-DAN that involve a multi-token prediction strategy. Interestingly, the attention weights of the MT-DAN even spread over the last predicted characters ("he " in the example) without disturbing the recognition. The combination of multi-token predictions and window queries enables the Meta-DAN to efficiently predict more tokens (9 here) by taking advantage of more context.
DAN, predicting "w"
MT-DAN ($m=5$), predicting "will "
W-DAN ($w=5$), predicting "will "
Meta-DAN ($m=5$, $w=5$), predicting "will stor"
Visualization of the attention process for a single decoding step. Attention weights are more diffused when relying on multi-token predictions (MT-DAN, Meta-DAN)
VI-E Large-scale comparison of the prediction approaches
CER results (in percentages) on the test sets with FCN-M encoder. We used $m=5$ and $w=5$ for the proposed approaches. Best results at page level are shown in bold. Dataset Line-level Page-level FCN DAN Faster DAN W-DAN (ours) MT-DAN (ours) Meta-DAN (ours) CER Time CER Time CER Time CER Time CER Time CER Time IAM 4.84 0.03 4.23 3.39 4.16 0.64 4.87 0.77 4.51 0.83 4.79 0.43 BRESSAY 3.45 0.02 2.43 24.65 2.34 1.51 3.02 6.63 2.33 5.70 2.38 4.01 READ-2016 4.61 0.03 3.85 4.88 4.34 0.73 4.34 1.09 3.80 0.80 4.22 0.63 Esposalles 0.61 0.05 5.86 19.99 1.91 2.23 5.95 4.91 2.16 3.37 2.13 2.86 ScribbleLens 4.72 0.06 4.38 16.24 4.81 2.17 5.06 4.04 4.16 2.71 4.31 2.24 CASIA-2 2.91 0.05 1.10 2.81 3.17 0.57 2.81 0.69 1.32 0.52 2.01 0.43 Eparchos 7.79 0.04 7.27 9.14 7.52 0.97 8.18 2.11 8.14 1.52 7.36 1.21 RIMES-2009 6.73 5.87 6.56 0.98 6.26 1.44 4.94 1.09 4.64 0.83 RIMES-2024 3.89 5.89 3.57 0.94 3.48 1.38 2.97 1.06 3.26 0.80 MAURDOR-C3 10.38 5.34 9.76 0.78 8.05 1.27 13.71 1.16 7.93 0.69 MAURDOR-C4 8.28 6.70 10.46 1.32 6.93 1.61 6.62 1.20 6.59 0.92 Average 5.31 9.54 5.33 1.17 5.36 2.36 4.97 1.81 4.51 1.37 # parameters 17.7 M 23.1 M 23.1 M 23.1 M 23.4 M 23.4 M
We now provide a large evaluation at page level over all datasets and all proposed prediction approaches, as well as results for the concurrent Faster DAN decoding strategy. We used the same architecture and training strategy for all the evaluating settings, i.e., we also train the concurrent approaches with the same framework for fair comparisons. We also provide results at line level, which corresponds to the pre-training stage (same encoder with CTC decoding) for comparison. However, let recall that line-level metrics only correspond to recognition on humanly-segmented images: it does not include the required segmentation and ordering stages that would lead to higher CER and prediction times. Results shown in Table VI-E highlight that it is generally beneficial to process documents at page level rather than line level. A counter-example is the Esposalles dataset, where line recognition performs very well, demonstrating a certain ease of visual reading (single writer). In this particular case, it seems that the benefits of implicit language modeling provided by the transformer decoder are not sufficient to compensate for the added difficulty of learning the reading order.
Another point is about the prediction time: the three proposed approaches are on par with the prediction speed of the Faster DAN. However, one can note that the Faster DAN speed improvement is highly dependent to the number of lines to be predicted in the document, contrary to the proposed approaches. Indeed, the Faster DAN is 16.3 times faster than the DAN for the BRESSAY dataset but 4.9 times faster for the CASIA-2 dataset, given that there are on average 30 lines per document for BRESSAY and 10 lines per document for CASIA-2.
Regarding the CER performance now, the MT-DAN approach reaches new state-of-the art performance over 5 datasets. Also, it seems that the enlarged context provided by the Meta-DAN enables to stabilize the attention process, reducing the phenomenon of predictions loops (when the same text is predicted multiple times) for heterogeneous layouts as in the MAURDOC-C3 and MAURDOR-C4 datasets. It results in new state-of-the-art recognition performance when averaging on all datasets with an improvement of 0.46 point compared to MT-DAN and 0.80 point compared to DAN. WER results are given in Appendix C for comparison purposes.
Comparison with state-of-the-art approaches (no external data, no external language model). Best result at page level is shown in bold. State-of-the-art line-level results are given for comparison. [b] Dataset Approaches of the literature CER (%) for our approaches (page level) Architecture Level CER (%) MT-DAN Meta-DAN BRESSAY [42] CRNN + attention line 3.62 2.33 2.38 [42] YOLO-v5 + Calamari line 6.01 [42] DANIEL page 3.55 CASIA-2 [48] CNN + attention line 2.34 1.32 2.01 [49] DenseNet + attention line 1.78 Eparchos [44] CNN + GRU line 4.53 8.14 7.36 Esposalles [21] DAN + NER line 0.48 2.16 2.13 [21] DAN + NER page 3.61 IAM [50] transformer line 3.99 4.51 4.79 [13] DAN page* 4.54 [17] DANIEL page* 4.38 MAURDOR-C3 [13] DAN page 8.62 13.71 7.93 [14] Faster DAN page 8.93 MAURDOR-C4 [13] DAN page 8.02 6.62 6.59 [14] Faster DAN page 9.88 READ-2016 [12] VAN line 4.10 3.80 4.22 [13] DAN paragraph 3.22 [13] DAN page 3.43 [14] Faster DAN page 3.95 [17] DANIEL page 4.03 RIMES-2009 [13] DAN page 4.54 4.94 4.64 [14] Faster DAN page 6.38 [17] DANIEL page 5.80 RIMES-2024 2.97 3.26 ScribbleLens [51] CNN + BLSTM line 6.47 4.16 4.31 * For IAM, the page level refers in the literature to pages for which the header including duplicated printed text was removed.
VI-F Comparison with the state of the art
We further compare the proposed approaches to the state of the art in Table VI-E. For fair comparison, we only compare to works following the same constraints, i.e., without external language model and that do not perform training on additional external dataset. Among the 7 datasets for which there is a page-level evaluation in the literature, we reach new state-of-the-art CER on 4 of these datasets with the MT-DAN or the Meta-DAN. The proposed approaches also outperform state-of-the-art line-level recognition models on ScribbleLens and CASIA-2 datasets. On can note an exception with the Eparchos dataset, which follow a similar behavior compared to the Esposalles dataset, i.e., text recognition is better at line level than at page level. This can be explained by the very few available training data that prevent from fully benefiting from the full potential of attention-based recognition systems. Lastly, we provide a first evaluation for the newly released RIMES-2024 dataset with a CER of 2.97% reached by the MT-DAN.
VI-G Multilingual recognition
CER (in percentages) for different encoders on the test sets for line-level multilingual recognition. FCN-M ConvNeXtV2-N ConvNeXtV2-B Pre-training IN-21k IN-21k IN-21k Training Monolingual Multilingual Monolingual Multilingual Multilingual Multilingual IAM 4.84 6.40 5.54 9.57 6.34 5.31 BRESSAY 3.45 5.27 3.61 7.85 4.51 3.23 READ-2016 4.61 6.92 5.23 8.28 6.25 4.74 Esposalles 0.61 1.83 1.02 2.11 1.33 0.82 ScribbleLens 4.72 6.49 5.27 8.61 6.07 5.07 Average CER (%) 3.65 5.38 4.13 7.28 4.90 3.83 # parameters 17.6 M 16.0 M 88.8 M Average time (s) 0.03 0.01 0.03
In this last experiment, we propose training the Meta-DAN in a multilingual setting, i.e., the model is trained on a set of datasets at once, and evaluated on the same datasets. We only focus on latin languages to see if learning one language can be helpful for learning another one when sharing part of the same alphabet.
We compare the FCN-M encoder we used so far to two variants of the ConvNeXt-V2 architecture [52], which is a state-of-the-art CNN architecture for computer vision, pre-trained through self-supervised learning (masked auto-encoder) on ImageNet-21k (IN-21k). Regarding these ConvNeXt-V2 encoders, we made the following modifications for fair comparison and CTC requirements for pre-training (minimum input length with respect to the ground truth): 1) the last projection is set to output 256 feature maps to be compatible with the DAN decoder; 2) The ConvNeXt encoder includes five downsampling stages, each one dividing by two the horizontal and vertical axes. We removed two of them for the horizontal axis, leading to a final output shape for the image features of ($H_{f}=\frac{H}{32},W_{f}=\frac{W}{8},C_{f}=256$), which corresponds to the FCN-M encoder output shape.
We first concentrate on the line-level pre-training stage in Table VI-G. As one can note, the FCN-M encoder is better suited for monolingual setting compared to the same-scale ConvNeXt-V2-N model, but we can observe the contrary for the multilingual setting. The results highlight that is mainly due to the self-supervised pre-training strategy on a large-scale IN-21k dataset: the overall CER is 2.38 point higher when training from scratch. Another interesting point is that, even with a larger model (ConvNeXt-V2-B, 4 times more parameters than FCN-M), the obtained results are still a bit less good as the monolingual approach with FCN-M encoder (3.83% of CER compared to 3.65%).
We now compare these different encoders at page level with the Meta-DAN approach. To make training more efficient, we add an extra training step at line-level, leading to a 4-step training: 1) pre-training the encoder with CTC supervision through the line-level approach; 2) transfer learning to MT-DAN on line-level images; 3) further training the MT-DAN at page level with curriculum learning 4) transfer learning to Meta-DAN at page level without curriculum strategy.
Table VI-G describes the obtained results, highlighting the need for an upscaled encoder to get close to monolingual results. Indeed, the ConvNeXt-V2-B encoder leads to an average CER of 4.62% on a multilingual setting, which is very close to the 4.45% obtained by the FCN-M with monolingual training. In addition, it only results in a small prediction time increase, from 1.57 second per page (FCN-M, monolingual) to 1.74 second per page (CN2-B, multilingual).
Comparison of CER results (in percentages) at page-level with FCN-M and ConvNeXt-V2 (CN2) encoders for multilingual training using the Meta-DAN ($w=5$, $m=5$). Dataset Monolingual Multilingual FCN-M FCN-M CN2-N CN2-B BRESSAY 2.38 4.82 4.14 4.07 READ-2016 4.22 5.43 6.80 4.91 Esposalles 2.13 2.58 2.69 2.47 ScribbleLens 4.31 9.79 7.51 5.87 IAM 4.79 5.10 4.89 3.97 RIMES-2024 3.26 3.22 3.31 2.93 MAURDOR-C3 7.93 7.41 6.91 7.84 MAURDOR-C4 6.59 17.94 7.06 4.92 Average CER 4.45 7.04 5.41 4.62 Average time (s) 1.57 1.36 1.01 1.74 # parameters 23.2 M 23.2 M 21.5 M 94.3 M
These results also demonstrates that the multilingual setting is only beneficial for similar languages regarding the evaluated models. Comparing monolingual FCN-M and multilingual CN2-B, one can notice that CER of all French/English datasets (IAM, RIMES-2024 and MAURDOR) are improved, while all datasets of historical documents (Esposalles, READ-2016 and ScribbleLens), which involves a specific period and language, are degraded.
VII Discussion
Through this extensive study on prediction strategies, we showed the robustness of the proposed Meta-DAN approach through many evaluation configurations (various datasets, multilingual recognition), reaching on-average state-of-the-art results. However, one can note that there is no one strategy that is better than the others on all datasets. We noticed that, no matter the prediction variant (Faster DAN or ours), it seems to have a beneficial impact on the loop effects that can occur during decoding, noticeable through the CER results of the Esposalles dataset (Table VI-E). We assume that for the Faster DAN strategy, this is due to the parallelization of the line recognition. Indeed, the loops mainly occur when switching from one line to another. For the approaches we proposed, this must be the result of the increased context modeling and the forced anticipation of the model to predict the following characters, which can be found on two consecutive lines. But contrary to the Faster DAN, the Meta-DAN does not suffer from "catastrophic" errors. Indeed, any missed or duplicated character in the first decoding stage of the Faster DAN lead to discard or duplicate a whole text line. While rare, this can severely impact the recognition performance, also explaining the recognition gap. Although we limit ourselves to compare the approaches for character-level predictions to reach state-of-the-art results with few training data, it has to be noted that all approaches could also be used in combination with BPE or N-gram tokenization to further reduce the prediction time.
VIII Conclusion
In this paper, we proposed and evaluated two new prediction strategies for end-to-end full-page text recognition: windowed queries and multi-token predictions. We show that the Meta Document Attention Network approach, combining both strategies, reaches new on-average state-of-the-art results on many handwritten datasets while reaching competitive prediction times. In addition, this approach enables flexible balance between prediction time and recognition performance by adapting the associated hyperparameters to adapt to the use case. While we only evaluated the Meta-DAN for HTR, the proposed approach is generic enough to be adapted to any autoregressive transformer-based system. Lastly, we highlight that using a generic self-supervised pre-trained encoder is an effective way to bridge the performance gap between monolingual and multilingual settings.
Acknowledgment
This work was granted access to the HPC resources of IDRIS under the allocation 2024-AD011012155R2 made by GENCI. Experiments presented in this paper were also partly carried out using the Grid’5000 testbed, supported by a scientific interest group hosted by Inria and including CNRS, RENATER and several Universities as well as other organizations (see https://www.grid5000.fr).
APPENDIX
VIII-A Scaling up the DAN encoder
We noticed that the DAN lacked of some feature extraction capability when evaluated on challenging datasets such as BRESSAY which presents some crossed-out text with annotations. We opt for scaling up the DAN encoder only in order to increase the performance while preventing a significant prediction time increase. Indeed, the encoding part of the model is only used once at inference (it is excluded from the iterative decoding stage); so the computation involved are not such significant compared to decoding. The proposed encoder is referred to as FCN-M, and the original one as FCN-S.
The proposed modification mainly consists in doubling the number of feature maps per block while preserving the last one which condition the size of the decoder, as detailed in Table VIII-A. We use the notation of the original paper [12] where ConvBlock is a Convolution Block made up of three convolutions, instance normalization and dropout, and DSCBlock is equivalent except for the convolutions that are replaced by Depthwise Separable Convolutions [36].
Composition of the two versions of the encoder. ConvBlock(X) and DSCBlock(X) indicate the number of feature maps X returned by each convolutional layer. FCN-S [12] FCN-M (ours) Output size layers ConvBlock(16) ConvBlock(32) $H\times W$ ConvBlock(32) ConvBlock(64) $\frac{H}{2}\times\frac{W}{2}$ ConvBlock(64) ConvBlock(128) $\frac{H}{4}\times\frac{W}{4}$ ConvBlock(128) ConvBlock(256) $\frac{H}{8}\times\frac{W}{8}$ ConvBlock(128) ConvBlock(512) $\frac{H}{16}\times\frac{W}{8}$ ConvBlock(128) ConvBlock(512) $\frac{H}{32}\times\frac{W}{8}$ DSCBlock(128) DSCBlock(512) $\frac{H}{32}\times\frac{W}{8}$ DSCBlock(128) DSCBlock(512) $\frac{H}{32}\times\frac{W}{8}$ DSCBlock(128) DSCBlock(512) $\frac{H}{32}\times\frac{W}{8}$ DSCBlock(256) DSCBlock(256) $\frac{H}{32}\times\frac{W}{8}$ params 1.7 M 17.6 M
We evaluate these modifications on READ-2016, CASIA-2, BRESSAY and IAM datasets at line level, and at page level with the DAN. Results are shown in Table VIII-A. As one can note, the modified encoder improves the recognition results (or is equivalent) on all the evaluated datasets, whether it is at line or page level. The prediction time is only slightly impacted, we assume this is mainly due to the complexity of learning the reading order. For CASIA-2 and IAM, the prediction time is lower, which can be explained by a reducing of the loop effect (predicting several times the same text part). But for READ-2016 and BRESSAY, the prediction time is higher, which can be the result of the additional computations due to the bigger encoder, and to a better recognition (missed text parts are now well recognized, adding extra decoding iterations). The most negative impact is on the READ-2016 dataset with a 6% increase of the prediction time. These results also enhance what have been found in [13]: the page-level DAN outperforms its line-level counterpart while they are based on the same encoder.
Impact of the encoder changes at line (pre-training stage) and page levels. CER are expressed in percentages, and times in seconds. Dataset Encoder FCN (line) DAN (page) CER Time CER Time READ-2016 FCN-S 4.82 0.03 3.82 4.61 FCN-M 4.61 0.03 3.85 4.88 CASIA-2 FCN-S 5.33 0.05 1.35 2.94 FCN-M 2.91 0.05 1.10 2.81 Bressay FCN-S 6.11 0.02 4.53 24.28 FCN-M 3.45 0.02 2.43 24.65 IAM FCN-S 5.76 0.03 4.53 3.57 FCN-M 4.84 0.03 4.23 3.39
VIII-B Datasets
Line-level splits
Table VIII-B details the split for training, validation and test used for the different datasets regarding the experiments at line level.
Datasets splits at line level. Dataset Train Validation Test IAM 6,482 976 2,915 READ-2016 8,367 1,043 1,140 Esposalles 2,328 742 757 ScribbleLens 4,408 375 563 CASIA-2 36,305 5,465 10,449 Eparchos 1,512 380 380 BRESSAY 19,561 4,608 5,915
Example images
Figure VIII-B provides document image examples from the 10 evaluated datasets. As one can note, these datasets cover many page layouts, visual aspects, writing styles, periods and languages.
MAURDOR-C3
MAURDOR-C4
IAM
CASIA-2
RIMES-2024
BRESSAY
ScribbleLens
Eparchos
Esposalles
READ-2016
Dataset samples from test sets. The MAURDOR-C3 and MAURDOR-C4 datasets provide widely heterogeneous documents in terms of layout, while the other datasets are quite homogeneous.
VIII-C WER results for the comparison of prediction strategies.
Table VIII-C provides the WER scores associated to experiments related to Section VI-E, i.e., when relying on monolingual evaluation and the FCN-M encoder. Interestingly, the WER results do not follow exactly the same trend as that of the CER, by exhibiting a slight advantage for the Faster DAN with reaches an on-average WER of 13.73%, compared to 14.22% for the Meta-DAN. There is no clear explanation for this observation. However, it still remains lower than the DAN WER which reaches 15.15%.
WER results (in percentages) on the test sets with FCN-M encoder. We used $m=5$ and $w=5$ for the proposed approaches. Best results at page level are shown in bold. Dataset Line-level Page-level FCN DAN Faster DAN W-DAN MT-DAN Meta-DAN IAM 15.83 14.00 13.42 16.28 14.96 15.78 BRESSAY 12.12 7.85 8.03 9.72 8.28 8.46 READ-2016 20.41 14.69 14.83 16.90 14.90 16.03 Esposalles 2.61 6.60 2.73 6.63 3.09 3.09 ScribbleLens 20.38 19.66 18.64 21.19 18.92 19.71 Eparchos 26.29 22.36 22.36 24.71 23.39 23.69 RIMES-2009 14.35 12.99 15.12 12.67 12.96 RIMES-2024 8.72 8.30 9.35 8.00 8.72 MAURDOR-C3 28.43 19.82 20.71 27.46 20.23 MAURDOR-C4 14.80 16.19 13.50 13.15 13.48 Average 15.15 13.73 15.41 14.48 14.22

### v2

@@
 However, the naïve character-level autoregressive decoding process results in long prediction times: it requires several seconds to process a single page image on a modern GPU.
-We propose the Meta Document Attention Network (Meta-DAN) as a novel decoding strategy to reduce the prediction time while enabling a better context modeling.
-It relies on two main components: windowed queries, to process several transformer queries altogether, enlarging the context modeling with near future; and multi-token predictions, whose goal is to predict several tokens per query instead of only the next one.
+We propose the Meta Document Attention Network (Meta-DAN) as a novel decoding strategy to reduce the prediction time while enabling better context modeling.
+It relies on two main components: windowed queries, to process several transformer queries altogether, enlarging the context modeling with the near future; and multi-token predictions, whose goal is to predict several tokens per query instead of only the next one.
 We evaluate the proposed approach on 10 full-page handwritten datasets and demonstrate state-of-the-art results on average in terms of character error rate.
 Source code and weights of trained models are available at https://github.com/FactoDeepLearning/meta_dan.
-Index Terms:
+keywords:
 Handwritten Text Recognition , multi-token prediction , Document Attention Network , full-page.
-I Introduction
+journal: Pattern Recognition
+[…]
+organization=Univ Rennes, CNRS, IRISA - UMR 6074, F-35000 Rennes,country=France
+1 Introduction
 Handwritten Text Recognition (HTR) is the process of automatically extracting handwritten text from a digitized image and converting it into machine-readable text.
 The historical approach to tackle this task, still in the majority today, consists of a multi-stage paradigm.
-A document image is segmented into sections of text, which are then ordered so as to preserve a coherent reading order.
-Thereafter, these text sections are recognized and reassembled to obtain the overall transcription of the document.
-Over time, the nature of the textual section has broadened, from the isolated character [1, 2], to the word [3, 4], then to the line [5, 6, 7] and more recently to the paragraph [8, 9, 10, 11, 12], gradually increasing the context usable during recognition.
-However, regardless of the granularity used, this multi-step approach comes with significant drawbacks: 1) the segmentation step relies on a neural network that needs to be trained, requiring additional segmentation annotations, 2) the ordering step is generally a hand-crafted algorithm that needs to be adapted to the layout of each target dataset, 3) the recognition of the individual sections is carried out independently, limiting the context modeling during prediction, 4) errors accumulate in each of these steps.
-Also, it has to be noted that works of the literature generally focus on the segmentation or recognition task separately, so that the given recognition results are based on a human-generated segmentation, which is not representative of the task in real-life conditions.
+A document image is segmented into sections of text, which are then ordered to preserve a coherent reading order.
+Thereafter, these text sections are recognized and reassembled to obtain the overall transcription.
+Over time, the nature of the textual section has broadened, from the isolated character [1], to the word [21], then to the line [45, 32, 12] and more recently to the paragraph [4, 3, 48, 15], gradually increasing the context usable during recognition.
+However, regardless of the granularity, this multi-step approach comes with significant drawbacks: 1) the segmentation step requires an additional segmentation model and annotations, 2) the ordering step is generally a dataset-specific hand-crafted algorithm, 3) the recognition of the individual sections is carried out independently, limiting the context modeling during prediction, 4) errors accumulate in each of these steps.
 More recently, the Document Attention Network (DAN) [14] was proposed as the first end-to-end approach for HTR at page level, achieving similar recognition performance compared to the multi-step approach.
-In addition to text, the DAN also predicts special XML-like layout tokens, leading to enriched serialized representations of documents.
 The DAN architecture includes a transformer-based decoder to perform a character-level autoregressive prediction process.
-Indeed, the transformer’s attention mechanism is used to generate an implicit segmentation of characters, one after the other, making the model learn the reading order of the document through a text-only supervision.
 Due to its state-of-the-art performance, the DAN is currently the default approach used for end-to-end handwritten text recognition at page-level [16, 8].
-In addition, it has been extended to handle other tasks at page level such as named entity recognition [16, 17, 18], table recognition [19], and key-value information extraction [20, 21, 22], by adding dedicated tokens in the predictable token set.
-The main drawback of the DAN, which is also an advantage, lies in its character-level iterative prediction process, which is slow when processing whole pages.
-Indeed, on the one hand, processing text at character level reduces the vocabulary used, making training more easy.
-It also makes it possible to manipulate lower-dimensional latent states, since each token carries little information, reducing computational complexity.
-On the other hand, it increases the size of the generated sequences, and thus, the number of iterations to be performed and the prediction time.
+In addition, it has been extended to handle other tasks at page level such as named entity recognition [11, 10], table recognition [5], and key-value information extraction [41, 40], by adding dedicated tokens in the predictable token set.
+The main drawback of the DAN, which is also an advantage, lies in its character-level iterative prediction process.
+Indeed, on the one hand, processing text at character level reduces the vocabulary, making training easier.
+It also enables to manipulate lower-dimensional latent states, since each token carries little information, reducing computational complexity.
+On the other hand, it increases the target sequence length, and so the number of decoding iterations, increasing the prediction time.
 In this work, we propose the Meta Document Attention Network (Meta-DAN), a novel prediction strategy to improve both inference speed and recognition accuracy for end-to-end page-level handwritten text recognition.
@@
 The first one is based on multi-token predictions as it showed interesting results for Natural Language Processing (NLP) [23].
-The second consists in windowed queries, i.e., processing successive transformer queries altogether as a sliding window to process them in parallel.
-The main contributions of our work are summarized as follows:
+The second consists of windowed queries, i.e., processing successive transformer queries altogether as a sliding window to process them in parallel.
+Our main contributions are as follows:
   • We propose two prediction strategies to improve page-level handwritten text recognition: windowed queries and multi-token predictions.
-  • We study the introduction of a specific dynamic prediction policy to strike a balance between prediction speed and recognition performance for multi-token predictions.
-  • We provide an extensive evaluation over a set of 10 datasets, covering several languages and alphabets.
-It highlights that combining both proposed approaches (Meta-DAN) leads to on-average improved recognition performance compared to concurrent approaches.
-  • We show that multilingual page-level text recognition can be beneficial to increase the recognition performance for the mainly represented languages.
-II Related Works
-II-A Handwritten text recognition
-Over the last decades, the field of handwriting recognition has made major advances, moving from isolated character recognition based on expert models, to end-to-end recognition of full-page documents.
-These advances are mainly due to the modeling capability of deep neural networks, which represent the state of the art in this field.
-The approaches proposed in the literature can be divided into two main categories we refer to as the alignment-based and the attention-based approaches.
-The alignment-based approaches consider the recognition task as a one-dimensional sequence alignment problem between a prediction sequence and a ground truth sequence.
-Modeling the text recognition task in this way enables the end-to-end recognition of word [24, 25] and line [5, 7] images.
-The practical application of this method is based on two aspects: the reduction of the vertical dimension, generally achieved through convolution or pooling, and the use of the Connectionist Temporal Classification (CTC) [26] introduced in 2006.
-The CTC loss enables to align a probability lattice computed from the input image with a sequence of characters corresponding to the ground truth.
-The strength of this approach is also its weakness: it only deals with one-dimensional sequences.
-To go a step further, two works showed that it is feasible to recognize whole paragraphs by using tricks to assimilate a paragraph to a long line of text, through bilinear interpolations [10], or concatenation [27].
-However, to the best of our knowledge, this alignment-based approach is not suited for end-to-end full-page recognition, since the complexity of page layouts (e.g., multi-column pages) makes it impossible to reduce it to a one-dimensional sequence alignment problem.
+  • We study the introduction of a confidence-based dynamic prediction policy to strike a balance between prediction speed and recognition performance for multi-token predictions.
+  • We provide an extensive evaluation over 10 various public datasets.
+To our knowledge, this is the largest benchmark evaluated at page-level.
+This evaluation highlights that combining both proposed approaches (Meta-DAN) leads to on-average state-of-the-art recognition performance.
+2 Related Works
+2.1 Handwritten text recognition
+State-of-the-art approaches can be divided into two main categories we refer to as the alignment-based and the attention-based approaches.
+The alignment-based approaches consider the recognition task as a one-dimensional sequence alignment problem between a predicted probability lattice and a ground truth sequence.
+The Connectionist Temporal Classification (CTC) [24] was proposed in 2006 to perform this alignment, enabling the end-to-end recognition of word [25, 7] and line [45, 12] images.
+However, the CTC only deals with one-dimensional sequences.
+This way, tricks have been proposed to assimilate a paragraph to a long line of text, through bilinear interpolations [48], or concatenation [13].
+But, to our knowledge, alignment-based approaches are not suited for end-to-end full-page recognition, since the complexity of page layouts (e.g., multi-column pages) prevent to reduce it to a one-dimensional sequence alignment problem.
 More recently, the attention-based approaches emerged, particularly with the rise in popularity of the Transformer architecture [43].
-They rely on attention mechanisms to provide an iterative decoding process at character level, performing a kind of implicit character segmentation through the attention maps.
-In other words, characters are predicted one after the other following an autoregressive process.
-At each decoding step, attention weights are computed and matched to the extracted features of the input image, with the aim of focusing on a specific part of the image (the character to recognize).
-Contrary to the alignment-based approaches, the attention-based approaches can be applied to any input image, from a single line [6, 29] or paragraph [9, 30, 11] to a two-page document with the DAN-based works [13, 14, 17].
-The main drawback of such approach lies in the prediction time that grows with the length of the text sequence to be recognized.
+They rely on attention mechanisms to provide an iterative autoregressive decoding process at character level, performing a kind of implicit character segmentation through the attention maps.
+Contrary to the alignment-based approaches, the attention-based approaches can be applied to any input image, from a single line [32, 2] or paragraph [3, 38] to a two-page document with the DAN-based works [14, 16, 11].
 It has to be noted that marginal works [4, 15] were proposed to handle paragraph recognition with a hybrid approach combining the attention mechanisms, to generate an implicit line segmentation, and the alignment-based approach, to recognize them iteratively through CTC decoding.
-In addition, an attempt of adapting the CTC to 2D-to-1D alignment was proposed in [31], but it did not lead to convincing results.
-In summary, alignment-based approaches are advantageous because of their one-shot prediction, but they are limited to images with a single column of text, whereas attention-based approaches are capable of handling entire pages with complex layout but involve longer prediction times.
-Attention-based decoding strategies also suffer from a limited context modeling, in that it only has access to the previously predicted tokens to make its prediction (only past is considered).
-In this work, we focus on reducing the prediction time of the attention-based character-level approach to reach fast end-to-end page-level predictions, while improving the context modeling through the integration of near future in the decoding process.
-II-B Speeding up autoregressive processes
-The prediction time issue of autoregressive processes can be found in various applications including a text generation stage, and several strategies have been proposed.
-In [32], a transformer-based network is trained to predict the $n$ future tokens instead of a single one in the context of text summarization.
-We refer to this strategy as multi-token prediction.
-This idea was successfully applied for large language models recently [23].
-In this latter work, $k$ prediction heads are applied on the same transformer output, each one predicting one of the $k$ next tokens, limiting the computational overhead.
-A similar approach was proposed for neural machine translation in [33] by including an extra verification step: prediction is performed another time using all the predictions as in teacher forcing to validate them or not.
-The authors of [34] proposed another way to reduce the prediction time for the task of image captioning.
-The target sequence is split into subgroups of words, which are decoded in parallel.
-They use an extra DEL token to remove duplicated predictions.
+In addition, an attempt to adapt the CTC to 2D-to-1D alignment was proposed in [36], but it did not lead to convincing results.
+In summary, alignment-based approaches are advantageous because of their one-shot prediction, but they are limited to images with a single column of text, whereas attention-based approaches are capable of handling entire pages with complex layouts, but involve longer prediction times due to their autoregressive nature.
+In this work, we focus on reducing the prediction time of the attention-based character-level approach to reach fast end-to-end page-level predictions, while improving the context modeling through the integration of the near future in the decoding process.
+2.2 Speeding up autoregressive processes
+The prediction time issue of autoregressive decoders is not limited to HTR.
+In [33], a multi-token prediction strategy was proposed to speed up text summarization: a transformer model is trained to predict the $n$ future tokens at once using $n$ projection heads.
+This idea was also successfully applied for neural machine translation [39] and for large language models [23].
+The authors of [22] proposed another way to reduce the prediction time for the task of image captioning. The target sequence is split into subgroups of words, which are decoded in parallel.
 Closer to our work, a similar approach was proposed to speed up the Document Attention Network by splitting the decoding process into two stages [16].
 The first stage is dedicated to predicting the first character of each line, and the second one consists in completing all lines in parallel, reducing the number of decoding steps.
-While the prediction time is significantly decreased, this alternative suffers from a quite chaotic context: the model has access to the beginning of each line, i.e., a partial view of both past and future.
 The windowed queries we propose also rely on the idea of processing queries in parallel, but they differ in that the processed queries are successive, leading to a unique decoding stage and a context mixing full past and near future.
-An alternative to speed up the prediction process is to reduce the length of the target sequence.
-This can be achieved through a specific tokenizer, one of the best known is the Byte-Pair Encoding (BPE) algorithm [35].
-Instead of encoding the text to be recognized as a sequence of character tokens, it is encoded as tokens representing sub-words of variable sizes, obtained from their frequency of occurrence in a given dataset.
-This technique has been applied in [17] showing a significant prediction time decrease, in return for a higher character error rate.
-The approach we propose is compatible with this kind of technique as it rely on the token prediction strategy, regardless of the nature of the tokens.
+An alternative to speed up the prediction process is to reduce the length of the target sequence, by relying on a dedicated tokenizer, such as the Byte-Pair Encoding (BPE) algorithm [37].
+This technique was applied in [11], showing a significant prediction time decrease, in return for a higher character error rate.
+The approach we propose is compatible with this kind of technique as it relies on the token prediction strategy, regardless of the nature of the tokens.
 In this paper, we propose to combine both windowed queries and multi-token predictions to speed up the prediction process while improving the context modeling capability of the model.
-While evaluated for the HTR task in this paper, the approach is generic and could be applied for any task relying on a transformer-based iterative decoding process.
-III Background
+3 Background
+3.1 HTR problem formulation
+[…]
+3.2 Document Attention Network
+The Document Attention Network (DAN) [14] was proposed to tackle this task as a character-level autoregressive prediction process.
+It refers to a deep neural network illustrated in Figure 3.2.
+It is made up of two main building blocks: a Fully Convolutional Network (FCN) encoder and a causal transformer-based decoder.
+The general idea of this approach is as follows.
+Based on the image features extracted by the encoder, and on the previously predicted tokens, the goal is to predict the next token (i.e., a character).
+To determine when to stop this iterative decoding process, a special end-of-transcription token (<e>) is added to the set of predictable tokens $\mathcal{A}$, and appended to the ground truth sequence: $\boldsymbol{y}\in\mathcal{A}^{L^{y}+1}$ with $\mathcal{A}=\mathcal{A}^{c}\cup\{\text{<e>}\}$ and $\boldsymbol{y}_{L^{y}+1}=\text{<e>}$.
 DAN architecture.
 Encoding stage: 2D features $\boldsymbol{f}$ are extracted from the input image through a Fully Convolutional Network (FCN).
-Features are enhanced with 2D positional encoding and flattened for transformer needs.
-Decoding stage: at each iteration $t$, a query sequence $\boldsymbol{q}^{t}$ is generated from the previous predictions $\hat{\boldsymbol{y}}^{t}$ and associated to the image features through transformer’s attention mechanisms to compute the new prediction $\hat{\boldsymbol{y}}_{t}$.
-To make things clearer, we first propose a formulation of the handwritten text recognition problem, we go over the main points of the DAN [13] approach on which is built this work, and we discuss the Faster DAN [14] proposition.
-III-A Problem formulation
-[…]
-III-B Document Attention Network
-The Document Attention Network (DAN) [13] was proposed to tackle this task as a character-level autoregressive prediction process.
-While the original approach generates additional layout tokens in addition to text tokens, we only focus on text recognition in this work.
-The DAN refers to a deep neural network illustrated in Figure III.
-It is made up of two main building blocks: a Fully Convolutional Network (FCN) encoder and a causal transformed-based decoder.
-The general idea of the approach is as follows.
-Based on the image features extracted by the encoder, and on the previously predicted tokens, the goal is to predict the next token (each token representing a character).
-The DAN encoder, noted $e$, was first proposed in [12] as a generic feature extractor for handwritten text recognition able to deal with images of variable sizes.
-It mainly consists in a succession of 18 standard convolutions and 12 Depthwise Separable Convolutions [36], coupled with instance normalization, mix dropout layers and skip connections.
-It extracts 2D features $\boldsymbol{f}=e(\boldsymbol{X})$ ($\boldsymbol{f}\in\mathbb{R}^{H_{f}\times W_{f}\times C_{f}}$) from the input image by reducing the height by 32 and the width by 8 ($H_{f}=\frac{H}{32}$, $W_{f}=\frac{W}{8}$ and $C_{f}=256)$.
+Decoding stage: at each iteration $t$, a query sequence $\boldsymbol{q}^{t}$ is generated from the previous predictions $\hat{\boldsymbol{y}}^{t}$ and associated with the image features through the transformer’s attention mechanisms to predict the next token $\hat{\boldsymbol{y}}_{t}$.
+The decoding is initialized with a start-of-prediction token $\hat{\boldsymbol{y}}_{0}=\text{<s>}$ and stops when the end-of-transcription token <e> is predicted.
+Regarding the notations, given that matrices change over the decoding iterations, we opted for the following convention.
+The exponent specifies the version of the matrix at a given time, and the index specifies a vector from that matrix.
+For instance, in Figure 3.2, $\boldsymbol{q}^{t}$ is the matrix of query token vectors at decoding iteration $t$ made up of $t$ vectors indexed from 0 to t1t-1.
+Given that the vector $\boldsymbol{q}_{i}^{t}$ at index $i$ is the same for all $t$, it can be referred to as $\boldsymbol{q}_{i}$, omitting the exponent for readability.
+In the following, we describe the different stages implied by the DAN by referring to the circled numbers indicated in Figure 3.2.
+1
+The DAN encoder, noted $\text{Enc}_{\text{DAN}}$, extracts 2D features $\boldsymbol{f}=\text{Enc}_{\text{DAN}}(\boldsymbol{X})$ ($\boldsymbol{f}\in\mathbb{R}^{H_{f}\times W_{f}\times C_{f}}$) from the input image by reducing the height by 32 and the width by 8 $(H_{f}=\frac{H}{32}$, $W_{f}=\frac{W}{8}$ and $C_{f}=256)$.
+This encoder was first proposed in [15] as a generic feature extractor for handwritten text recognition able to deal with images of variable sizes.
 All details can be found in the original paper [15].
-[…]
-For each iteration $t$, the previously predicted tokens are put together as query input for the decoder: $\hat{\boldsymbol{y}}^{t}=[\hat{\boldsymbol{y}}_{0},\dots,\hat{\boldsymbol{y}}_{t-1}]$.
-A trainable weight matrix is used to embed each text token as a vector of dimension $C_{f}$, leading to $\boldsymbol{e}\in\mathbb{R}^{t\times C_{f}}$.
-Following the original transformer [28], one-dimensional positional encoding ($\boldsymbol{p}^{t}_{\text{1D}}$) is added, leading to the query sequence $\boldsymbol{q}^{t}=\boldsymbol{e}^{t}+\boldsymbol{p}^{t}_{\text{1D}}$.
-In the same way, its 2D adaptation [30] encoding both dimensions of the image on half of the channels is used to inject positional information to the image features, resulting in $\boldsymbol{g}=\boldsymbol{f}+\boldsymbol{p}_{\text{2D}}$.
-Query tokens $\boldsymbol{q}^{t}$ and images features $\boldsymbol{g}$ are processed through 8 stacked transformer decoder layers resulting in $\boldsymbol{o}^{t}=d(\boldsymbol{q}^{t},\boldsymbol{g})$, preserving the number of channels $C_{f}$.
-Finally, a score $\boldsymbol{s}_{t}$ for each token of the token set $\mathcal{A}$ is computed through a linear projection layer $p$ based on the output corresponding to the last token of the query sequence:
-$\boldsymbol{s}_{t}=p(\boldsymbol{o}_{t-1}).$ (1)
-The predicted token corresponds to the one with the highest score: $\hat{\boldsymbol{y}}_{t}=\operatorname*{arg\,max}(\boldsymbol{s}_{t})$.
+2 For transformer needs, positional information is injected into the image features $\boldsymbol{f}$ through additive 2D positional encoding $\boldsymbol{\pi}_{\text{2D}}$ [38].
+Height and width are each encoded over half of the vector channels.
+The result is reshaped to a one-dimensional representation: $\boldsymbol{g}=\text{flatten}(\boldsymbol{f}+\boldsymbol{\pi}_{\text{2D}})$.
+3 For each iteration $t$, the previously predicted tokens are put together to be used by the model to know which character token should be predicted next: $\hat{\boldsymbol{y}}^{t}=[\hat{\boldsymbol{y}}_{0},\dots,\hat{\boldsymbol{y}}_{t-1}]$.
+A special start-of-transcription token (<s>) is used to initiate the prediction process: $\hat{\boldsymbol{y}}_{0}=\text{<s>}$.
+4 A trainable weight matrix $\boldsymbol{E}\in\mathbb{R}^{(|\mathcal{A}|+1)\times C_{f}}$ is used to map each token of the vocabulary $\mathcal{A}$ to a vector of dimension $C_{f}$, leading to $\boldsymbol{e}^{t}\in\mathbb{R}^{t\times C_{f}}$, with $\boldsymbol{e}^{t}_{i}=\boldsymbol{E}_{\hat{\boldsymbol{y}}_{i}}$.
+One can note the additional embedding vector (+1), which corresponds to the start-of-transcription token (<s>).
+5 As the transformer attention mechanism is permutation equivariant, one-dimensional positional encoding $\boldsymbol{\pi}^{t}_{\text{1D}}$ [43] is added to these embeddings to preserve the order information of the tokens, resulting in the query sequence $\boldsymbol{q}^{t}=\boldsymbol{e}^{t}+\boldsymbol{\pi}^{t}_{\text{1D}}$.
+6 The query tokens $\boldsymbol{q}^{t}$ are processed through the DAN decoder, noted $\text{Dec}_{\text{DAN}}$, which corresponds to 8 stacked transformer decoder layers preserving the number of channels $C_{f}$.
+The goal is to extract relevant information from the flattened image features $\boldsymbol{g}$, i.e., a latent representation $\boldsymbol{o}^{t}_{t-1}$ of the next character to predict: $\boldsymbol{o}^{t}=\text{Dec}_{\text{DAN}}(\boldsymbol{q}^{t},\boldsymbol{g})$.
+7 Based on the last output of the decoder $\boldsymbol{o}_{t-1}$, a score vector $\boldsymbol{s}_{t}\in\mathbb{R}^{|\mathcal{A}|}$ is computed through a linear projection layer of trainable weights $\boldsymbol{W}\in\mathbb{R}^{|\mathcal{A}|\times C_{f}}$ to associate a score with each predictable token:
+$\boldsymbol{s}_{t}=\boldsymbol{W}\boldsymbol{o}_{t-1}.$ (1)
+8
+The predicted token $\hat{\boldsymbol{y}}_{t}$ corresponds to the one with the highest score: $\hat{\boldsymbol{y}}_{t}=\operatorname*{arg\,max}(\boldsymbol{s}_{t})$.
+Steps 3 to 8 are repeated until the specific end-of-transcription token (<e>) is predicted.
 The DAN is trained using the cross-entropy loss $\mathcal{L}_{\text{CE}}$ over the sequence of logits of the scores obtained through softmax activation:
-$\mathcal{L}_{\text{DAN}}=\sum_{t=1}^{L+1}\mathcal{L}_{\text{CE}}(\boldsymbol{y}_{t},\boldsymbol{p}_{t}),$ (2)
-with $\boldsymbol{p}_{t,i}=\operatorname*{soft\,max}(\boldsymbol{s}_{t})=\frac{e^{\boldsymbol{s}_{t,i}}}{\sum_{j}e^{\boldsymbol{s}_{t,j}}}$.
-[…]
-[width=]images/tikz/dan_in_out DAN [13]
-[…]
-[width=]images/tikz/fdan_in_out Faster DAN [14]
-[…]
-[width=]images/tikz/wdan_in_out W-DAN (ours, $w=3$)
-[…]
-[width=]images/tikz/mt-dan_in_out MT-DAN (ours, $m=3$)
-[…]
-[width=0.928]images/tikz/meta_dan_in_out Meta-DAN (ours, $w=3$, $m=2$)
-[…]
-III-C Faster DAN
-The Faster DAN [13] was proposed as an alternative to speed up the decoding process at inference stage.
-It has to be noted that the Faster DAN architecture is exactly the same as the DAN.
-The contribution is about decoding strategy whose aim is to parallelize the text line predictions.
-It relies on a two-stage approach.
+$\mathcal{L}_{\text{DAN}}=\sum_{t=1}^{L^{y}+1}\mathcal{L}_{\text{CE}}(\boldsymbol{y}_{t},\boldsymbol{p}_{t}),$ (2)
+where $\boldsymbol{p}_{t,i}=\frac{e^{\boldsymbol{s}_{t,i}}}{\sum_{j}e^{\boldsymbol{s}_{t,j}}}$ is the predicted probability associated with token $i$ at time $t$.
+3.3 Faster DAN
+The Faster DAN [16] was proposed as a decoding alternative to speed up the inference stage.
+It aims to parallelize the text line predictions by relying on a two-stage approach.
 First, an autoregressive process predicts the first character of each line, following the original DAN strategy.
@@
 In this way, for $N$ text lines, there are $N$ token predictions per iteration.
-Equation 1 becomes:
-$\boldsymbol{s}^{l}_{t}=p(\boldsymbol{o}^{l}_{t-1}),$ (3)
-where $\boldsymbol{s}^{l}_{t}$ is the score vector associated to the token prediction of the next character for line $l$ at iteration $t$, and $\boldsymbol{o}^{l}_{t-1}$ is the transformer output corresponding to the last predicted token for line $l$.
-The Faster DAN introduces a specific positional encoding to associate each text token with a line index and a positioning index within that line.
-This is a crucial component to make this parallelization possible.
+Equation 1 becomes: $\boldsymbol{s}^{l}_{t}=\boldsymbol{W}\boldsymbol{o}^{l}_{t-1}$, where $\boldsymbol{s}^{l}_{t}$ is the score vector associated with the token prediction of the next character for line $l$ at iteration $t$, and $\boldsymbol{o}^{l}_{t-1}$ is the transformer output corresponding to the last predicted token for line $l$.
 Although the Faster DAN showed interesting inference speed improvements, it results in a loss of past context when predicting a character.
 Indeed, instead of benefiting from the whole past context as is the case for the DAN, the Faster DAN has a partial view of the past and future due to the parallel line predictions, which reduces its ability to model the language.
-In this paper, we propose a strategy to get the best of both worlds: a reduced prediction time through query parallelization, and an improved context modeling by processing successive queries altogether and relying on multi-token predictions.
-IV Meta Document Attention Network
+The strategy we propose aims to get the best of both worlds: improving context modeling by processing successive queries altogether through multi-token predictions, while reducing prediction time through parallelization.
+4 Meta Document Attention Network
 We propose two new approaches for the autoregressive prediction of text tokens in the context of end-to-end handwritten text recognition at page level.
@@
 We assume this is a key element to reduce the number of recognition errors through language modeling.
-First, we propose to associate multiple predictions to a single query $q_{t-1}$, to predict the $m$ next tokens at once.
-This is referred as the Document Attention Network with Multi-Tokens predictions (MT-DAN).
+First, we propose to perform multiple predictions from the same query $\boldsymbol{q}_{t-1}$, to predict the $m$ next tokens at once.
+This is referred to as the Document Attention Network with Multi-Tokens predictions (MT-DAN).
 Second, we propose a query approach based on a window mechanism to process multiple queries at once, each associated with a single prediction.
-This is referred as the Document Attention Network with Windowed queries (W-DAN).
+This is referred to as the Document Attention Network with Windowed queries (W-DAN).
 Finally, we design a global strategy including these two approaches, called Meta Document Attention Network (Meta-DAN).
-A visual comparison of the prediction processes is depicted in Figure III-B.
-IV-A MT-DAN
+A visual comparison of the prediction processes is depicted in Figure 4.
+[…]
+[width=]dan_in_out DAN [14]
+[…]
+[width=]fdan_in_out Faster DAN [16]
+[…]
+[width=]wdan_in_out W-DAN ($w=3$)
+[…]
+[width=]mt-dan_in_out MT-DAN ($m=3$)
+[…]
+[width=0.85]meta_dan_in_out Meta-DAN ($w=3$, $m=2$)
+[…]
+4.1 MT-DAN
 Inspired by recent works on multi-token predictions for Natural Language Processing (NLP) [23], we propose the Document Attention Network with Multi-Tokens predictions (MT-DAN).
@@
 Equation 1 becomes:
-$\boldsymbol{s}^{(\boldsymbol{t})}_{t+k}=p^{k}(\boldsymbol{o}_{t-1}),$ (4)
-where $\boldsymbol{s}^{(\boldsymbol{t})}_{t+k}$ is the score vector associated to the token prediction of the $k^{\text{th}}$ next token computed at iteration $t$, $p^{k}$ is the $k^{\text{th}}$ projection head, and $\boldsymbol{o}_{t-1}$ is the transformer output corresponding to the last query token $\boldsymbol{q}_{t-1}$.
+$\boldsymbol{s}^{(\boldsymbol{t-1})}_{t+k}=\boldsymbol{W}^{k}\boldsymbol{o}_{t-1},$ (3)
+where $\boldsymbol{s}^{(\boldsymbol{t-1})}_{t+k}$ is the score vector associated with the token prediction of the $k^{\text{th}}$ next token computed at iteration $t$.
+The exponent $(\boldsymbol{t-1})$ indicates that this prediction was performed based on query indexed by t1t-1.
+$\boldsymbol{W}^{k}$ is the weight matrix corresponding to the $k^{\text{th}}$ projection head, and $\boldsymbol{o}_{t-1}$ is the transformer output corresponding to the last query token $\boldsymbol{q}_{t-1}$.
 In this way, $m$ predictions are performed per iteration following:
@@
 with $k\in[0,\dots,m-1]$.
-As one can note, several predictions $\hat{\boldsymbol{y}}^{(\boldsymbol{t})}_{t+k}$ will be associated to the same target token $\boldsymbol{y}_{t+k}$.
-This is not an issue during training, we train the MT-DAN using the cross-entropy loss over the whole token sequence, as for the original DAN.
-It involves adding $m-1$ extra <e> tokens in order for all text tokens to have a target for each of the projection heads, leading to the following loss:
-$\mathcal{L}_{\text{MT-DAN{}}}=\sum_{t=1}^{L+1}\sum_{k=0}^{m-1}\mathcal{L}_{\text{CE}}(\boldsymbol{y}_{t+k},\boldsymbol{p}^{(\boldsymbol{t})}_{t+k}),$ (6)
+As one can note, several predictions $\hat{\boldsymbol{y}}^{(\boldsymbol{t})}_{t+k}$ are associated with the same target token $\boldsymbol{y}_{t+k}$.
+This is not an issue during training; we train the MT-DAN using the cross-entropy loss over the whole token sequence, as for the original DAN.
+It involves adding m1m-1 extra <e> tokens in order for all text tokens to have a target for each of the projection heads, leading to the following loss:
+$\mathcal{L}_{\text{MT-DAN{}}}=\sum_{t=1}^{L^{y}+1}\sum_{k=0}^{m-1}\mathcal{L}_{\text{CE}}(\boldsymbol{y}_{t+k},\boldsymbol{p}^{(\boldsymbol{t})}_{t+k})$, (5)
 with $\boldsymbol{p}^{(\boldsymbol{t})}_{t+k}=\operatorname*{soft\,max}(\boldsymbol{s}^{(\boldsymbol{t})}_{t+k})$.
-In Section VI, we propose a decoding policy based on the confidence level of the prediction to choose not to predict all $m$ next tokens for poor confidence.
-We demonstrate it is a good trade-off between prediction speed and recognition performance.
 Predicting multiple tokens $\hat{\boldsymbol{y}}_{t+k}$ from a unique output $\boldsymbol{o}_{t-1}$ can theoretically lead to saturation as $m$ increases.
 Indeed, the fixed-size vector $\boldsymbol{o}_{t-1}$ must represent the next $m$ text tokens at once.
-To solve this issue, we propose another way to predict multiple tokens per iteration by processing several consecutive queries at once.
-IV-B W-DAN
-[…]
-[width=]images/tikz/dan_attention DAN [13]/MT-DAN (ours)
-[…]
-[width=]images/tikz/fdan_attention Faster DAN [14]
-[…]
-[width=]images/tikz/wdan_attention W-DAN (ours) / Meta-DAN (ours)
-(for $w=3$)
-[…]
+To solve this issue, we propose another way to predict multiple tokens per iteration, which consists in processing several consecutive queries at once.
+4.2 W-DAN
 To solve the MT-DAN output compression issue, we propose the Document Attention Network with Windowed queries (W-DAN).
 Instead of relying on a single output to predict $m$ tokens, we opt for using the output of the last $w$ predictions to predict the $w$ next tokens, $w$ being a new hyperparameter.
-In this way, it solves the compression issue as they are as many output as they are predictions.
-More precisely, each output $\boldsymbol{o}_{t-q}$ is responsible for the prediction of token $\hat{\boldsymbol{y}}_{t-q+w}$, with $q\in[1,\dots,w]$.
+In this way, it solves the compression issue as there are as many outputs as there are predictions.
+The decoding process of the W-DAN acts as a sliding window over the query tokens, processing them by chunks of $w$ tokens at once, leading to $w$ new outputs.
+More precisely, each output $\boldsymbol{o}_{t-q}$ is responsible for the prediction of a single token $\hat{\boldsymbol{y}}_{t-q+w}$, with $q\in[1,\dots,w]$.
 Equation 1 becomes:
-$\boldsymbol{s}_{t-q+w}=p(\boldsymbol{o}_{t-q}),$ (7)
+$\boldsymbol{s}_{t-q+w}=\boldsymbol{W}\boldsymbol{o}_{t-q},$ (6)
 with
 $\hat{\boldsymbol{y}}_{t-q+w}=\operatorname*{arg\,max}(\boldsymbol{s}_{t-q+w}).$ (7)
-[…]
-$\mathcal{L}_{\text{W-DAN{}}}=\sum_{t=w}^{L+1+n_{\text{e}}}\mathcal{L}_{\text{CE}}(\boldsymbol{y}_{t},\boldsymbol{p}_{t}),$ (9)
-There is a major advantage of both presented approaches compared to the DAN and the Faster DAN: the $k$ or $w$ predictions follow one another and are performed altogether.
-It means that the context is enlarged when prediction takes place.
+It has to be noted that all scores $\boldsymbol{s}_{t-q+w}$ are computed using the same projection head of weights $\boldsymbol{W}$ as they are applied independently on each $\boldsymbol{o}_{t-q}$.
+In addition, it is necessary to initialize the decoding process with $w$ initial tokens to predict the first $w$ tokens.
+We arbitrarily choose to use the same start-of-transcription token $w$ times ($\hat{\boldsymbol{y}}_{0}=\cdots=\hat{\boldsymbol{y}}_{w-1}=\text{<s>}$).
+We also added $n_{\text{e}}$ extra end-of-transcription tokens at the end of the target sequence so as to make the target length divisible by $w$.
+It leads to the following loss:
+$\mathcal{L}_{\text{W-DAN{}}}=\sum_{t=w}^{L^{y}+1+n_{\text{e}}}\mathcal{L}_{\text{CE}}(\boldsymbol{y}_{t},\boldsymbol{p}_{t}).$ (8)
+The major advantage of both presented approaches compared to the DAN and the Faster DAN is that the $k$ or $w$ predictions are consecutive and performed altogether: the context is enlarged when the prediction takes place.
 The model must focus on a broader context to predict multiple successive tokens.
 For instance, for $k=3$ (or $w=3$), the model focuses on predicting $\{\hat{\boldsymbol{y}}_{t},\hat{\boldsymbol{y}}_{t+1},\hat{\boldsymbol{y}}_{t+2}\}$ at once.
-This way, the prediction of $\hat{\boldsymbol{y}}_{t}$ is impacted by the future context $\{\hat{\boldsymbol{y}}_{t+1},\hat{\boldsymbol{y}}_{t+2}\}$, which improves the ability to model the language at prediction time compared with DAN (which only benefits from past context).
-The same goes with $\hat{\boldsymbol{y}}_{t+1}$ predicted with $\{\hat{\boldsymbol{y}}_{t+2}\}$ as context of future.
-IV-C Meta-DAN
-We propose now to formulate a generic prediction strategy to take advantage of both presented approaches by combining them.
-It is referred as the Meta Document Attention Network (Meta-DAN).
+This way, the prediction of $\hat{\boldsymbol{y}}_{t}$ is impacted by the future context $\{\hat{\boldsymbol{y}}_{t+1},\hat{\boldsymbol{y}}_{t+2}\}$, which improves the ability to model the language at prediction time compared to DAN.
+4.3 Meta-DAN
+We propose to take advantage of both presented approaches by combining them: this is referred to as the Meta Document Attention Network (Meta-DAN).
 Thus, the Meta-DAN is parametrized by two hyperparameters $m$ and $w$, indicating the number of projection heads and the window size, respectively.
@@
 $\boldsymbol{o}_{t-q}$, is responsible for the prediction of $\{\hat{\boldsymbol{y}}_{t-q+w},\dots,\hat{\boldsymbol{y}}_{t-q+w+m-1}\}$ as follows:
-$\hat{\boldsymbol{y}}^{(\boldsymbol{t})}_{t}=\operatorname*{arg\,max}(\boldsymbol{s}^{(\boldsymbol{t})}_{t}),$ (10)
+$\hat{\boldsymbol{y}}^{(\boldsymbol{t-q})}_{t-q+w+k}=\operatorname*{arg\,max}(\boldsymbol{s}^{(\boldsymbol{t-q})}_{t-q+w+k})$, (9)
 with
-$\boldsymbol{s}^{(\boldsymbol{t-q+1})}_{t-q+w+k}=p^{k}(\boldsymbol{o}_{t-q}),$ (11)
-where $\boldsymbol{s}^{t-q+1}_{t-q+w+k}$ is the score associated to token $\hat{\boldsymbol{y}}_{t-q+w+k}$ by the $k^{\text{th}}$ projection head of the $q^{\text{th}}$ query token in the query window, and with $k\in[0,\dots,m-1]$ and $q\in[1,\dots,w]$.
-The Meta-DAN involves initializing the decoding process with $w$ <s> tokens, as for the W-DAN, and adding extra <e> so that the last query has a target prediction for all its associated projection heads.
+$\boldsymbol{s}^{(\boldsymbol{t-q})}_{t-q+w+k}=\boldsymbol{W}^{k}\boldsymbol{o}_{t-q},$ (10)
+where $\boldsymbol{s}^{(\boldsymbol{t-q})}_{t-q+w+k}$ is the score associated with token $\hat{\boldsymbol{y}}_{t-q+w+k}$ by the $k^{\text{th}}$ projection head of the $q^{\text{th}}$ query token in the query window, and with $k\in[0,\dots,m-1]$ and $q\in[1,\dots,w]$.
+The Meta-DAN involves initializing the decoding process with $w$ <s> tokens, as for the W-DAN, and adding $n_{e}$ extra <e> so that the last query has a target prediction for all its associated projection heads.
 Regarding training, equation 2 becomes:
-$\mathcal{L}_{\text{Meta-DAN{}}}=\sum_{t=w}^{L+1+n_{\text{e}}}\sum_{k=0}^{m-1}\mathcal{L}_{\text{CE}}(\boldsymbol{y}_{t+k},\boldsymbol{p^{(\boldsymbol{t})}_{t+k}})$ (12)
-Detailed features of evaluated datasets (page-level).
-Dataset Train Val Test Charset Average Language Century Writers size # char.
-per page IAM (paragraph) [37] 747 161 336 80 371 English 20th-21st 657 RIMES-2009 [38] 1,050 100 100 108 588 French 21st $\approx 400$ RIMES-2024 [39] 4,500 505 600 123 567 French 21st $\approx$ 1,400 READ-2016 [40] 350 50 50 89 468 German 15th-19th unknown Esposalles [41] 75 25 25 62 1,748 Catalan 17th 1 BRESSAY [42] 647 154 199 109 2,612 Portuguese 21st 1,000 ScribbleLens [43] 164 14 21 85 1,410 Dutch 16th-18th 85 Eparchos [44] 80 20 20 157 962 Greek 16th 2 CASIA-2 [45] 3,600 475 1,015 2,704 273 Chinese 21st 1,019 MAURDOR-C3 [46] 1,006 148 166 134 491 French/English 21st > 130 MAURDOR-C4 [46] 721 111 114 127 675 French/English 21st > 90
+$\mathcal{L}_{\text{Meta-DAN{}}}=\sum_{t=w}^{L^{y}+1+n_{\text{e}}}\sum_{k=0}^{m-1}\mathcal{L}_{\text{CE}}(\boldsymbol{y}_{t+k},\boldsymbol{p}^{(\boldsymbol{t})}_{t+k})$ (11)
 As one can note, several query tokens can be in charge of the same prediction.
-This is shown in Figure III-B with a Meta-DAN example with $w=3$ and $m=2$: the "m" and "p" are predicted twice, from two different queries.
-In this work, we limit ourselves to use the prediction of the first projection head for the $w-1$ first queries, and the prediction of all projection heads for the last query.
-While discarded, we assume that training to predict those duplicated characters encourages the network to use its future context to improve language modeling.
-In this way, all queries follow the same behavior, which make training easier.
-In summary, the Meta-DAN approach is a general formulation for which one can distinguish three specific cases depending on the hyperparameters:
+This is shown in Figure 4 with an example with $w=3$ and $m=2$: the "m" and "p" are predicted twice, from two different queries.
+In this work, we limit ourselves to using the prediction of the first projection head for the w1w-1 first queries, and the prediction of all projection heads for the last one.
+Although discarded, we assume that training to predict these extra characters encourages the network to improve language modeling.
+In summary, the Meta-DAN is a general formulation for which one can distinguish three specific cases depending on the hyperparameters:
   • Choosing $m=1$ and $w=1$ consists in the DAN approach.
@@
   • Choosing $m=1$ and $w>1$ leads to the W-DAN approach.
-IV-D Impact on the computational complexity
-To further give insights about the main differences between all these prediction strategies, Figure IV-B provide visuals of the attending maps (transformer masks), stating which token can access to which other tokens.
-Regarding the MT-DAN approach, the attending maps of the DAN and the MT-DAN are identical; this means that they both imply the same computations in the transformer decoder layers.
-In addition, contrary to the NLP application which relies on the Byte-Pair Encoding algorithm leading to large vocabularies, the addition of extra projection heads do not significantly impact the global computational complexity as we are working at character level, limiting the dimensions of the projection weight matrices.
+4.4 Impact on the computational complexity
+To further give insights into the main differences between all these prediction strategies, Figure 4.4 provide visuals of the attending maps (transformer masks), stating which token can access to which other tokens for the masked self-attention.
+[…]
+[width=]dan_attention DAN [14]/MT-DAN (ours)
+[…]
+[width=]fdan_attention Faster DAN [16]
+[…]
+[width=]wdan_attention W-DAN / Meta-DAN
+(ours, for $w=3$)
+[…]
+Regarding the MT-DAN approach, the attending maps of the DAN and the MT-DAN are identical, no matter the value of $m$; this means that they both imply the same computations in the transformer decoder layers.
+In addition, contrary to the NLP application, which relies on the Byte-Pair Encoding algorithm leading to large vocabularies, the addition of extra projection heads does not significantly impact the global computational complexity as we are working at character level, limiting the dimensions of the projection weight matrices.
 Concerning the W-DAN approach (and thus the Meta-DAN), the attending maps show that there are more interactions between the tokens in the self-attention layers, resulting in more computations, theoretically.
-However, to our knowledge, the current implementations of the transformer relies on masks to "hide" the forbidden tokens (usually the future) after the query-key multiplication takes place.
-In this way, this does not impact the computation time in practice since computations are performed over all token pairs whether they are really used or not.
-No matter the approach (DAN, Faster DAN, MT-DAN, W-DAN, Meta-DAN), the prediction process can be parallelized at training time using teacher forcing, so there is no additional benefit for the training stage.
-As stated previously, the number of computations are not significantly impacted neither by the MT-DAN nor by the W-DAN.
+However, to our knowledge, the current implementations of the transformer rely on masks to "hide" the forbidden tokens (usually the future) after the query-key multiplication takes place.
+In this way, this does not impact the computation time in practice since computations are performed over all token pairs, whether they are really used or not.
+No matter the approach, the prediction process can be parallelized at training time using teacher forcing [46], i.e., by injecting the whole ground truth instead of the autoregressive previous predictions, so that all decoding iterations are performed at once.
+In this way, there is no extra benefit for the training stage.
 The major impact is positive and related to the time complexity at prediction time when parallelizing the computation of token predictions.
-V Experimental conditions
-V-A Datasets
-We evaluate the proposed approach on many public HTR datasets, namely IAM [37], READ-2016 [40], Esposalles [41], BRESSAY [42], ScribbleLens [43], Esparchos [44], CASIA-2 [45], MAURDOR [46], RIMES-2009 [38] and RIMES-2024 [39].
-They cover 8 languages, three alphabets (Latin, Greek and Chinese), 6 centuries (15th-21st).
-It enables us to have a robust evaluation due to the high diversity of the datasets.
-Indeed, as detailed in Table IV-C, they involve many writing configurations (from single writer to more than a thousand), various sizes of character set, as well as a variable amount of training data (from less than a hundred to few thousand pages).
-We used the standard split into training, validation and test sets when existing.
-To our knowledge, we are the first to use the recently released RIMES-2024 dataset, so we propose a new split, randomly selecting samples for training, validation and test by respecting the split provided at line level in 2011 [47].
+5 Experimental conditions
+5.1 Datasets
+We select 10 public datasets with high diversity to make our evaluation robust: IAM [31], READ-2016 [35], Esposalles [34], BRESSAY [17], ScribbleLens [19], Esparchos [42], CASIA-2 [30], MAURDOR [6], RIMES-2009 [27] and RIMES-2024 [28].
+As detailed in Table 5.1, they cover 8 languages, three alphabets (Latin, Greek, and Chinese), and 6 centuries.
+They also involve many writing configurations (from a single writer to more than a thousand), various sizes of character sets, as well as a variable amount of training data.
+Detailed features of evaluated datasets (page-level).
+Dataset Train Val Test Charset Average Language Century Writers size # char/page IAM (paragraph) [31] 747 161 336 80 371 English 20th-21st 657 RIMES-2009 [27] 1,050 100 100 108 588 French 21st $\approx 400$ RIMES-2024 [28] 4,500 505 600 123 567 French 21st $\approx$ 1,400 READ-2016 [35] 350 50 50 89 468 German 15th-19th unknown Esposalles [34] 75 25 25 62 1,748 Catalan 17th 1 BRESSAY [17] 647 154 199 109 2,612 Portuguese 21st 1,000 ScribbleLens [19] 164 14 21 85 1,410 Dutch 16th-18th 85 Eparchos [42] 80 20 20 157 962 Greek 16th 2 CASIA-2 [30] 3,600 475 1,015 2,704 273 Chinese 21st 1,019 MAURDOR-C3 [6] 1,006 148 166 134 491 French/English 21st > 130 MAURDOR-C4 [6] 721 111 114 127 675 French/English 21st > 90
+We used the standard split into training, validation, and test sets when existing.
+To our knowledge, we are the first to use the recently released RIMES-2024 dataset, so we propose a split, randomly sampling for training, validation and test by respecting the split provided at line level in 2011 [26].
 We discarded images from the "Q" category, which corresponds to forms, since they do not provide the whole transcript.
-For multilingual experiments, we discard the RIMES-2009 dataset from the evaluation setting as some test samples can be found in the RIMES-2024 training set.
 For ScribbleLens, we removed 10% of the training samples to create a validation set by randomly picking from each of its data sources.
-For Eparchos, we propose a split at page-level by taking 80% of the samples for training, 10% for validation and 10% for test.
-Appendix B provides the splits at line level, used for pre-training purposes, for the datasets providing line-level annotations.
-This section also includes some input document image examples.
-[b] Comparison of W-DAN recognition performance with respect to the window size $w$ on the IAM and BRESSAY test sets. Speed Factor (SF) is indicated as a comparison with the DAN baseline.
-Window IAM BRESSAY
-size ($w$) CER (%) WER (%) Time (s) SF CER (%) WER (%) Time (s) SF
-1 (DAN) 4.23 14.00 3.39 x1.0 2.43 7.85 24.65 x1.0
-2 4.67 15.14 1.94 x1.7 2.91 8.87 20.86 x1.2
-3 4.83 15.62 1.26 x2.7 3.05 9.24 12.25 x2.0
-4 4.75 15.71 0.97 x3.5 2.63 9.35 8.63 x2.8
-5 4.87 16.28 0.77 x4.4 3.02 9.72 6.63 x3.7
-10* 5.07 16.53 0.40 x8.5 2.66 9.53 2.95 x8.4
-25* 6.56 20.94 0.18 x18.8 4.37 15.21 1.13 x21.8
-    Training was performed over 2M and 3M training samples for $w=10$ and $w=25$, respectively (instead of 1M).
-V-B Training strategy
+For Eparchos, we propose a split at page level, by taking 80% of the samples for training, 10% for validation, and 10% for test.
+A provides the splits at line level, used for pre-training purposes, for the datasets providing line-level annotations.
+This section also includes document image examples.
+5.2 Training strategy
 If not stated otherwise, we follow the main guidelines described in [14]:
-  • We pre-train the encoder part of the architecture using a line-level training strategy.
+  • The encoder is pre-trained using a CTC-based line-level training strategy.
 An adaptive max-pooling layer is added on top of the encoder to collapse the vertical axis, leading to a one-dimensional sequence of features.
-Then classification layer is applied to generate a probability lattice.
-This pre-training is carried out using the CTC loss.
+Then, a classification layer is applied to generate a probability lattice.
 If line-level annotations are provided in the dataset, we use a ratio of 25% real / 75% synthetic images (100% synthetic otherwise).
-  • We randomly apply some data augmentation strategies with a probability of 90%.
-It includes resolution modification, color jittering, Gaussian noising, Gaussian blurring, elastic distortion, dilation, erosion and perspective transformation.
-  • During page-level training, we rely on a curriculum synthetic document generation strategy.
-The generated documents are more and more complex through the training, starting with single line images to documents including several dozens lines.
-The model is trained on both real and synthetic documents with the proportion of synthetic documents slowly decreasing from 90% to 10%.
-Following [17], we scraped the website 1001-fonts.com to get a large collection of fonts from which we randomly sample when rendering a synthetic text line.
-The text used for synthetic generation comes from the training set of the target dataset itself.
+  • Data augmentation is applied with a probability of 90%, including resolution change, color jittering, Gaussian noising, Gaussian blurring, elastic distortion, dilation, erosion, and perspective transformation.
+  • The page-level training relies on a curriculum synthetic document generation strategy.
+The generated images contain more and more text lines through training.
+The ratio of synthetic documents slowly decreases from 90% to 10%.
+Following [11], we rely on the 1001-fonts.com website to get a large collection of fonts from which we randomly sample when rendering a synthetic text line.
+The text used for synthetic generation is from the training set of the target dataset.
+  • Line-level pre-training is carried out using 5M training samples; page-level training (including curriculum) is performed over 1M samples.
 Training and pre-training are performed using the Adam optimizer with an initial learning rate of $10^{-4}$, and automatic mixed-precision.
-  • Line-level pre-training is carried out using 5M training samples, and page-level training (including curriculum) is performed over 1M samples.
-  • Input images are preprocessed to reach a resolution of 150 dpi by preserving the original aspect ratio, whether it is for pre-training, training or evaluation.
-  • We do not use any external data, external language model nor lexicon constraints.
-V-C Evaluation
-We use the standard Character Error Rate (CER) and the Word Error Rate (WER) to evaluate the text recognition performance.
-Both are computed with the Levenshtein distance between the ground truth and the predicted text, normalized by the length of the ground truth.
+  • Preprocessing: images are scaled to reach 150 dpi resolution. No post-processing: we do not use any external data or language model.
+5.3 Evaluation
+We compute the standard Character Error Rate (CER) and Word Error Rate (WER) to evaluate the text recognition performance.
+Both metrics are computed using the Levenshtein distance [29], which counts the number of deletion, addition, and substitution operations required to go from the prediction sequence to the ground truth sequence, normalized by the length of the ground truth sequence.
+While the CER considers the prediction and the ground truth as sequences of characters, the WER considers them as sequences of words.
+All results are given for the test sets.
 It has to be noted that page-level approaches are sensitive to reading order; it means that correct text recognized in a wrong order will be considered as incorrect.
 For each model, we compute the average prediction time for one test image when using a single A100 GPU and a mini-batch size of 1.
-To provide fully comparable experiments, notably between monolingual and multilingual experiments we present in the following, we set the same maximum number of decoding iterations for all datasets.
-It is set to 5,000 to cover the requirements of all evaluated dataset.
-It has to be noted that this choice results in recognition metrics that can highly be impacted by the prediction loops that can occur, i.e., when the model predicts again and again the same text line.
-VI Experiments
-In this section, we evaluate the proposed approaches (W-DAN, MT-DAN and Meta-DAN) on the IAM and BRESSAY datasets to select the best approach and hyperparameters.
-While relying on the original DAN architecture [13], we propose an enlarged version of the encoder we refer to as FCN-M as it shows better recognition performance (details in Appendix A).
-Then, we apply them on the other 9 datasets to validate the robustness of the approach, and compare it to the state of the art.
-Finally, we present the results for a multilingual recognition setting.
-VI-A W-DAN
-We evaluate the W-DAN approach for window sizes $w\in\{1,2,3,4,5,10,25\}$, with $w=1$ corresponds to the DAN setting.
-Results are shown in Table V-A.
-It highlights that it is possible to predict from 2 to 10 character tokens at once, while still being competitive in terms of CER and WER.
-It is especially true for the BRESSAY dataset for which $w=10$ leads to a small increase of only 0.23 point of CER, for an 8.4 speed factor.
-However, results degrade significantly for $w=25$ with a CER of 6.56% instead of 4.23% for IAM, and 4.37% instead of 2.43% for BRESSAY, even with three times more training samples.
+We set the maximum number of decoding iterations for all datasets to 5,000 to cover the requirements of all evaluated datasets.
+This way, we do not rely on dataset-specific properties, which is closer to a real-case scenario.
+It has to be noted that this choice results in recognition metrics that can be highly impacted by the prediction loops that can occur, i.e., when the model predicts again and again the same text line.
+6 Experiments
+In this section, we evaluate the proposed approaches (W-DAN, MT-DAN, and Meta-DAN) on the IAM and BRESSAY datasets to select the best approach and hyperparameters.
+Then, we apply them to the other 9 datasets to validate the robustness of the approach.
+While relying on the original DAN architecture [14], we propose an enlarged version of the encoder, which we refer to as FCN-M as it shows better recognition performance (details in B).
+6.1 W-DAN
+Table 6.1(a) provides the results of the W-DAN approach for several window sizes $w$.
+It highlights that it is possible to predict up to 10 character tokens at once, while being competitive in terms of CER.
+It is especially true for the BRESSAY dataset for which $w=10$ leads to a small increase of only 0.23 points of CER.
+However, results degrade significantly for $w=25$ with a CER of 6.56% instead of 4.23% for IAM, and 4.37% instead of 2.43% for BRESSAY.
 As one can note, it is possible to reduce the average prediction time per sample nearly linearly with respect to the window size.
-This is illustrated in Figure VI-A.
-Regarding the BRESSAY prediction time, it appears that the new implementation has impacted the prediction time for large text predictions, leading to a lag in the speed factor improvement.
-A main observation is that the model is harder and harder to train as $w$ increases, so we needed to increase the amount of training data for $w=10$ and $w=25$ up to 2M and 3M, respectively.
-Evolution of the prediction time with respect to the window size $w$ for the W-DAN and to the number of token heads $m$ for the MT-DAN on the test set of IAM and BRESSAY datasets.
-The prediction time dramatically decreases as $m$ or $w$ increases at first, and then less sharply as the computation load of the encoder becomes more and more significant on the overall inference process.
-VI-B MT-DAN
-[b] Comparison of MT-DAN recognition performance with respect to the number of token heads $m$ on the IAM and BRESSAY test sets.
-Speed Factor (SF) is indicated as a comparison with the DAN baseline.
-Number of IAM BRESSAY heads ($m$) CER (%) WER (%) Time (s) SF CER (%) WER (%) Time (s) SF 1 (DAN) 4.23 14.00 3.39 x1.0 2.43 7.85 24.65 x1.0 2 4.70 14.85 1.84 x1.8 2.47 8.54 13.02 x1.9 3 4.74 15.17 1.23 x2.8 2.46 8.60 8.85 x2.8 4 4.68 15.56 0.93 x3.6 2.70 8.91 6.64 x3.7 5 5.14 16.01 0.75 x4.5 2.45 8.72 5.28 x4.7 10* 5.06 16.29 0.39 x8.7 2.37 8.64 2.64 x9.3 25* 7.19 23.83 0.16 x21.2 3.15 11.28 1.11 x22.2
-    Training was performed over 2M training samples for $m=10$ and $m=25$. An additional character-level curriculum stage was also added for the early stages of the training.
-We evaluate the MT-DAN approach in a similar setting, i.e., for various numbers of token heads $m\in\{1,2,3,4,5,10,25\}$, with $m=1$ corresponds to the original DAN.
-The obtained results (Table VI-B) are quite similar to those of the W-DAN, with comparable recognition performance and prediction time.
+Regarding BRESSAY prediction times, it appears that the new implementation has impacted the prediction time for large text predictions, leading to a lag in the speed factor improvement.
+A main observation is that the model is harder to train as $w$ increases.
+We assume that this is due to the inherent complexity of the self-attention layer, which grows with the length of the query sequence.
+Indeed, the more query tokens are processed altogether, the greater the required effort to coordinate and focus on the right parts of the image features.
+In this way, we needed to increase the amount of training data to reach convergence for $w=10$ and $w=25$ up to 2M and 3M, respectively.
+Comparison of W-DAN and MT-DAN performance with respect to the window size $w$ and to the number of token heads $m$, respectively, on the IAM and BRESSAY test sets.
+[b]
+(a) W-DAN results.
+$w$ IAM BRESSAY
+CER (%) Time (s) CER (%) Time (s)
+1 (DAN) 4.23 3.39 2.43 24.65
+2 4.67 1.94 2.91 20.86
+3 4.83 1.26 3.05 12.25
+4 4.75 0.97 2.63 8.63
+5 4.87 0.77 3.02 6.63
+10* 5.07 0.40 2.66 2.95
+25* 6.56 0.18 4.37 1.13
+(b) MT-DAN results.
+$m$ IAM BRESSAY
+CER (%) Time (s) CER (%) Time (s)
+1 (DAN) 4.23 3.39 2.43 24.65
+2 4.70 1.84 2.47 13.02
+3 4.74 1.23 2.46 8.85
+4 4.68 0.93 2.70 6.64
+5 5.14 0.75 2.45 5.28
+10* 5.06 0.39 2.37 2.64
+25* 7.19 0.16 3.15 1.11
+    More training data was used for $w=10$, $m=10$, $w=25$, and $m=25$
+6.2 MT-DAN
+For fair comparison, we evaluate the MT-DAN approach for the same values of token heads $m\in\{1,2,3,4,5,10,25\}$, as both $m$ for MT-DAN and $w$ for W-DAN refer to the number of predicted tokens per decoding iteration.
+Here, $m=1$ corresponds to the original DAN.
+The obtained results shown in Table 6.1(b) are quite similar to those of the W-DAN, with comparable recognition performance and prediction time.
 Neither approach seems to stand out from the other.
-Even for the extreme case of $m=25$, it appears that the MT-DAN performs better for the BRESSAY dataset with a CER of 3.15% compared to 4.37% for the W-DAN (with $w=25$), but we observe the contrary for the IAM dataset with a CER of 7.19% for the MT-DAN compared to 6.56% for the W-DAN.
-As for the W-DAN, it required more training data for $m=10$ and $m=25$.
-We also relied on a curriculum strategy at character level before the line-level curriculum strategy takes place.
-We did not use it for the W-DAN as it was not effective; we assume this is due to the way the tokens interact for make its predictions in this approach.
+Even for the extreme case of $m=25$, it appears that the MT-DAN performs better for the BRESSAY dataset with a CER of 3.15% compared to 4.37% for the W-DAN (with $w=25$), but it is the contrary for the IAM dataset: 7.19% of CER compared to 6.56% for the W-DAN.
+As for the W-DAN, it required more training data for $m=10$ and $m=25$ to reach convergence.
+However, this is not for the same reason.
+We assume that, for the MT-DAN, this is due to the compression task that is harder as $m$ increases.
+Indeed, $m$ character representations must be compressed into a single fixed-size vector.
+We also relied on a curriculum strategy at character level before the line-level curriculum strategy.
+We did not use it for the W-DAN as it was not effective; we assume this is due to the way the tokens interact to make their predictions in this approach.
 Improving the prediction policy
-To enhance the capacity of the MT-DAN approach, we propose and evaluate two prediction policies at inference time:
-1) The static prediction strategy consists in always predicting the $k$ next tokens at each decoding iteration ($k$ must be lower or equal to the number of tokens heads $m$).
-Previous results (shown in Table VI-B) relied on this naïve prediction policy, setting $k=m$.
-2) The dynamic prediction strategy involves predicting between the next 1 and $m$ tokens, based on a confidence threshold $\tau$.
-This way, the number of predicted tokens must vary from one decoding iteration to another.
-The idea is to accept to slow down the decoding process when the model is unsure of its prediction, by assuming that it could be better at the next decoding iteration for which it will benefit from more context.
-The first token head is always used to avoid infinite loop; the other $m-1$ heads are then considered sequentially.
-If the confidence score (i.e., the maximum value after softmax operation) associated to the current head is greater or equal to $\tau$, then the prediction is preserved, and the next head is processed; the decoding iteration ends otherwise.
+To enhance the capacity of the MT-DAN approach, we propose and evaluate two variants, we called prediction policies, at inference time:
+1) The static prediction policy consists in always predicting the $k$ next tokens at each decoding iteration ($k\leq m$).
+Previous results, shown in Table 6.1(b), relied on this naïve prediction policy, setting $k=m$.
+2) The dynamic prediction policy involves predicting between $1$ and $m$ tokens at once, based on a confidence threshold $\tau$.
+The number of predicted tokens must vary from one decoding iteration to another.
+The idea is to predict fewer tokens when the model is unsure of its prediction, by assuming that it could be better at the next decoding iteration, for which it will benefit from more context.
+The first token head is always used to avoid an infinite loop; the other m1m-1 heads are then considered sequentially.
+If the confidence score (i.e., the maximum value after softmax operation) associated with the current head is greater than or equal to $\tau$, then the prediction is preserved, and the next head is processed; the decoding iteration ends otherwise.
+Accuracy-speed trade-off with respect to confidence threshold $\tau$ for dynamic decoding of the MT-DAN on IAM and BRESSAY.
 We first study the choice of the confidence threshold $\tau$ for the dynamic strategy.
-Figure VI shows the evolution of the CER with respect to $\tau$ for values between 0% and 100%, with a step of 5%.
-We compare it for different numbers of heads and for both IAM and BRESSAY datasets.
-It appears that the recognition performance can globally be improved as of $\tau=0.5$, except for the configurations $m=3$ and $m=4$ on the BRESSAY dataset whose behavior is less monotonous.
-One should note that this prediction policy can even reach better CER performance than the DAN (blue line) on the BRESSAY dataset, for $w=2$ and $w=5$.
-We consider $\tau=0.9$ as a good compromise over all the evaluated configurations and use it for the next experiments.
+Figure 6.2 shows the evolution of the CER and of the prediction time with respect to $\tau$ for values between 0 and 1, with a step of 0.1.
+A confidence threshold of $\tau=0$ means that the token predicted by each projection head is always kept (the fastest option): it corresponds to the static prediction policy.
+A confidence threshold of $\tau=1$ means that only the prediction of the first projection head is kept (the longest option): it results in the DAN prediction strategy.
+We compare the accuracy-speed trade-off for different numbers of heads on both IAM and BRESSAY datasets.
+We also show results for a single projection head as a baseline, leading to a CER plateau as the dynamic strategy is designed for multi-token predictions.
+It appears that the confidence threshold fulfills its purpose, as there is a trend toward a decrease in the CER as $\tau$ increases, for both datasets.
+In addition, the prediction time per sample is rather stable from $\tau=0$ to $\tau=0.9$.
+However, one can note a huge prediction time gap going from $\tau=0.9$ up to $\tau=1$, making the approach ineffective as it consists in single-token prediction, as the DAN.
+In this way, we consider $\tau=0.9$ as a good compromise over all the evaluated configurations and use it for the next experiments.
+We compare the two prediction policies, namely static and dynamic, in Table 6.2.
+For the static prediction policy, we evaluate the models for $k\in\{1,2,3,4,5\}$ when possible.
+Regarding the IAM dataset, one can observe a trend whereby the CER decreases as $k$ also decreases, with optimum performance reached for $k=1$, noting that decreasing $k$ highly increases the prediction time.
+This trend is less noticeable on the BRESSAY dataset with more fluctuating results.
+However, the proposed dynamic prediction policy with $\tau=0.9$ shows interesting results, which generally outperform the corresponding model using the static strategy with $k=m$, with a limited prediction time overhead.
+For instance, it enables to reduce the CER from 5.14% to 4.51% for the IAM dataset ($m=5$), while slightly increasing the prediction time from 0.75 seconds (Table 6.1(b)) to 0.83 seconds.
 Comparison of static and dynamic prediction policies of the MT-DAN.
-CER (%) are given for the test sets of IAM and BRESSAY datasets.
+CER (%) are given for IAM and BRESSAY datasets.
 Best results are shown in bold; results that are very close to the best (whose difference in values is less than 0.05%) are underlined.
-Number IAM BRESSAY of heads static dynamic static dynamic ($m$) $k=1$ $k=2$ $k=3$ $k=4$ $k=5$ $\tau=0.9$ time (s) $k=1$ $k=2$ $k=3$ $k=4$ $k=5$ $\tau=0.9$ time (s) 1 (DAN) 4.23 2.43 2 4.33 4.70 4.59 1.92 2.41 2.47 2.36 13.60 3 4.31 4.85 4.74 4.31 1.32 2.46 2.46 2.46 2.69 9.21 4 4.48 4.54 4.68 4.68 4.52 1.00 2.71 2.86 2.61 2.70 2.71 6.97 5 4.45 4.74 4.54 4.86 5.14 4.51 0.83 2.34 2.40 2.36 2.40 2.45 2.33 5.70 10 4.79 4.80 4.87 4.87 4.90 4.80 0.47 2.67 2.24 2.70 2.28 2.29 2.66 3.15 25 5.73 5.61 5.58 5.64 5.68 5.47 0.36 3.35 3.30 2.92 2.91 2.93 2.87 1.88
-Impact of the CER threshold for dynamic decoding of the MT-DAN on the test set of IAM and BRESSAY.
-We compare the two prediction policies, namely static and dynamic, in Table VI.
-For the static prediction strategy, we evaluate the models for $k\in\{1,2,3,4,5\}$ when possible.
-Regarding the IAM dataset, one can observe a trend whereby the character error rate decreases as $k$ also decreases, with optimum performance reached for $k=1$, noting that decreasing $k$ highly increases the prediction time as shown previously.
-This trend is less noticeable on the BRESSAY dataset with more fluctuating results.
-However, the proposed dynamic prediction policy with $\tau=0.9$ shows interesting results which generally outperforms the corresponding model using the static strategy with $k=m$, with a limited prediction time overhead.
-For instance, it enables to reduce the CER from 5.14% to 4.51% for the IAM dataset ($w=5$), while slightly increasing the prediction time from 0.75 second (Table VI-B) to 0.83 second.
-VI-C Meta-DAN
-In this section, we focus on $m=5$ and $w=5$ as it seems to be a good comprise between prediction time and recognition performance.
-Pre-training strategy
-Comparison of Meta-DAN ($m=5$, $w=5$) pre-training strategy on the test sets of IAM and BRESSAY.
-Pre-training curriculum IAM BRESSAY line-level (baseline) 5.59 3.01 W-DAN ($w=5$) 4.82 2.51 W-DAN ($w=5$) 5.01 2.42 MT-DAN ($m=5$) 5.04 2.36 MT-DAN ($m=5$) 4.79 2.38
-In Table VI, we evaluate the Meta-DAN for several pre-training strategies.
-We compare the standard line-level pre-training to transfer learning directly from page-level W-DAN or MT-DAN, with and without curriculum learning.
-We observe that best results are obtained through page-level pre-training, highlighting the usefulness of the pre-trained decoder even if it does not rely on the same decoding strategy.
-We also note that re-training from line-level through curriculum learning is generally not necessary to reach the best results.
-Lastly, it seems that the MT-DAN pre-training leads to slightly better CER on average.
-For these reasons, the Meta-DAN training will be based on MT-DAN pre-training without curriculum learning in the following experiments.
+Number IAM BRESSAY of heads static dynamic static dynamic ($m$) $k=1$ $k=2$ $k=3$ $k=4$ $k=5$ $\tau=0.9$ Time $k=1$ $k=2$ $k=3$ $k=4$ $k=5$ $\tau=0.9$ Time 1 (DAN) 4.23 2.43 2 4.33 4.70 4.59 1.92 2.41 2.47 2.36 13.60 3 4.31 4.85 4.74 4.31 1.32 2.46 2.46 2.46 2.69 9.21 4 4.48 4.54 4.68 4.68 4.52 1.00 2.71 2.86 2.61 2.70 2.71 6.97 5 4.45 4.74 4.54 4.86 5.14 4.51 0.83 2.34 2.40 2.36 2.40 2.45 2.33 5.70 10 4.79 4.80 4.87 4.87 4.90 4.80 0.47 2.67 2.24 2.70 2.28 2.29 2.66 3.15 25 5.73 5.61 5.58 5.64 5.68 5.47 0.36 3.35 3.30 2.92 2.91 2.93 2.87 1.88
+6.3 Meta-DAN
+In this section, we focus on $m=5$ and $w=5$ as it seems to be a good compromise between prediction time and recognition performance.
+Through experiments, we found out that transfer learning from MT-DAN to Meta-DAN leads to better results than from W-DAN or line-level pre-training on both IAM and BRESSAY datasets, so we keep this setting in the following.
 Comparison with MT-DAN and W-DAN
-We compare the Meta-DAN with the MT-DAN and the W-DAN for the same hyperparameters ($m=5$ and $w=5$) in Table VI.
+We compare the Meta-DAN with the MT-DAN and the W-DAN for the same hyperparameters ($m=5$ and $w=5$) in Table 6.3.
+However, it should be noted that this is not a constraint of the proposition, and that $m$ and $w$ can be chosen completely independently for the Meta-DAN.
 As one can note for the IAM and BRESSAY datasets, combining both windowed queries and multi-token predictions enable to increase the recognition performance, as well as reducing the prediction time.
-Comparing to DAN, the conclusion is not as clear-cut regarding the character error rate as it is in favor to the DAN for IAM with a CER of 4.23% compared to 4.79%, but in favor of the Meta-DAN for BRESSAY with a CER of 2.38% compared to 2.43%.
-It shows that the Meta-DAN can be both more efficient and faster than the DAN depending on the target dataset.
-Comparison of proposed approaches using the static prediction policy (with $m=k$).
-CER are expressed in percentages, and times in seconds.
+Compared to DAN, the conclusion is not as clear-cut regarding the character error rate as it is in favor of the DAN for IAM with a CER of 4.23% compared to 4.79%, but in favor of the Meta-DAN for BRESSAY with a CER of 2.38% compared to 2.43%.
+It shows that the Meta-DAN can be both more efficient and faster than the DAN, depending on the target dataset.
+Comparison of proposed approaches using the static prediction policy ($m=k$).
 Approach $m$ $w$ IAM BRESSAY CER Time CER Time DAN 1 1 4.23 3.39 2.43 24.65 W-DAN 1 5 4.87 0.77 3.02 6.63 MT-DAN 5 1 5.14 0.75 2.45 5.28 Meta-DAN 5 5 4.79 0.43 2.38 4.01
-Comparing prediction strategies
-As the Meta-DAN also relies on multi-token predictions, we can apply the same dynamic prediction policy as for the MT-DAN.
-Its usage is evaluated in Table VI.
-It appears that the confidence scores associated to the predictions are not as high as for the MT-DAN, leading to a failure case for the BRESSAY dataset for which the prediction time is close to that of W-DAN (static case with $k=1$).
-As noted previously, the static policy with $k=1$ leads to the best recognition performance but does not benefit from accelerated prediction.
+Comparison of Meta-DAN ($w=5$, $m=5$) performance with respect to the decoding strategy.
+Strategy for IAM BRESSAY concurrent CER Time CER Time static ($k=1$) 4.55 0.79 2.31 7.16 static ($k=5$) 4.79 0.43 2.38 4.01 dynamic ($\tau=0.9$) 4.73 0.45 2.37 7.14
+Comparing prediction strategies.
+As the Meta-DAN relies on multi-token predictions, we can apply the same dynamic prediction policy as for the MT-DAN.
+This is evaluated in Table 6.3.
+It appears that the confidence scores associated with the predictions are not as high as those for the MT-DAN, leading to a failure case for the BRESSAY dataset for which the prediction time is close to that of W-DAN (static case with $k=1$).
+As previously, the static policy with $k=1$ leads to the best recognition performance but suffers from longer prediction.
 We consider the static policy with $k=m=5$ as the best alternative for the Meta-DAN and use this configuration in the following.
-Comparison of Meta-DAN ($w=5$, $m=5$) performance with respect to the decoding strategy on the test sets of IAM and BRESSAY.
-Strategy for IAM BRESSAY concurrent CER (%) Time (s) CER (%) Time (s) static ($k=1$) 4.55 0.79 2.31 7.16 static ($k=5$) 4.79 0.43 2.38 4.01 dynamic ($\tau=0.9$) 4.73 0.45 2.37 7.14
-VI-D Visualization
-A visualization of the attention process for the different prediction approaches is provided in Figure VI-D.
-Displayed images are cropped to a single line for visibility, but predictions are performed on the full-page images.
-Attention weights of the last cross-attention layer are projected on the input image and values are represented by color intensity (maximum intensity for 0.2 value).
-One color is associated to each projection head in this order: red, blue, green, yellow and cyan.
-As one can notice, the attention weights are more peaky for DAN and W-DAN.
-It was expected as each query only focuses on a single character prediction, contrary to MT-DAN and Meta-DAN that involve a multi-token prediction strategy.
-Interestingly, the attention weights of the MT-DAN even spread over the last predicted characters ("he " in the example) without disturbing the recognition.
+6.4 Visualization
+A visualization of the attention process for the different approaches is provided in Figure 6.4.
+Images are cropped to a single line for visibility, but predictions are performed on the full-page images.
+Attention weights of the last cross-attention layer are projected on the input image (values are normalized such that the maximum intensity is reached for 0.2).
+One color is associated with each projection head in this order: red, blue, green, yellow, and cyan.
+As expected, the attention weights are peakier for DAN and W-DAN since each query only focuses on a single character prediction, contrary to MT-DAN and Meta-DAN that involve a multi-token prediction strategy.
+Interestingly, the attention weights of the MT-DAN even spread over the last predicted characters ("he ") without disturbing the recognition.
 The combination of multi-token predictions and window queries enables the Meta-DAN to efficiently predict more tokens (9 here) by taking advantage of more context.
@@
 Visualization of the attention process for a single decoding step. Attention weights are more diffused when relying on multi-token predictions (MT-DAN, Meta-DAN)
-VI-E Large-scale comparison of the prediction approaches
+6.5 Large-scale comparison of the prediction approaches
 CER results (in percentages) on the test sets with FCN-M encoder.
 We used $m=5$ and $w=5$ for the proposed approaches.
-Best results at page level are shown in bold.
-Dataset Line-level Page-level FCN DAN Faster DAN W-DAN (ours) MT-DAN (ours) Meta-DAN (ours) CER Time CER Time CER Time CER Time CER Time CER Time IAM 4.84 0.03 4.23 3.39 4.16 0.64 4.87 0.77 4.51 0.83 4.79 0.43 BRESSAY 3.45 0.02 2.43 24.65 2.34 1.51 3.02 6.63 2.33 5.70 2.38 4.01 READ-2016 4.61 0.03 3.85 4.88 4.34 0.73 4.34 1.09 3.80 0.80 4.22 0.63 Esposalles 0.61 0.05 5.86 19.99 1.91 2.23 5.95 4.91 2.16 3.37 2.13 2.86 ScribbleLens 4.72 0.06 4.38 16.24 4.81 2.17 5.06 4.04 4.16 2.71 4.31 2.24 CASIA-2 2.91 0.05 1.10 2.81 3.17 0.57 2.81 0.69 1.32 0.52 2.01 0.43 Eparchos 7.79 0.04 7.27 9.14 7.52 0.97 8.18 2.11 8.14 1.52 7.36 1.21 RIMES-2009 6.73 5.87 6.56 0.98 6.26 1.44 4.94 1.09 4.64 0.83 RIMES-2024 3.89 5.89 3.57 0.94 3.48 1.38 2.97 1.06 3.26 0.80 MAURDOR-C3 10.38 5.34 9.76 0.78 8.05 1.27 13.71 1.16 7.93 0.69 MAURDOR-C4 8.28 6.70 10.46 1.32 6.93 1.61 6.62 1.20 6.59 0.92 Average 5.31 9.54 5.33 1.17 5.36 2.36 4.97 1.81 4.51 1.37 # parameters 17.7 M 23.1 M 23.1 M 23.1 M 23.4 M 23.4 M
-We now provide a large evaluation at page level over all datasets and all proposed prediction approaches, as well as results for the concurrent Faster DAN decoding strategy.
-We used the same architecture and training strategy for all the evaluating settings, i.e., we also train the concurrent approaches with the same framework for fair comparisons.
+Best results at page level are in bold.
+Dataset Line-level Page-level FCN DAN Faster DAN W-DAN MT-DAN Meta-DAN CER Time CER Time CER Time CER Time CER Time CER Time IAM 4.84 0.03 4.23 3.39 4.16 0.64 4.87 0.77 4.51 0.83 4.79 0.43 BRESSAY 3.45 0.02 2.43 24.65 2.34 1.51 3.02 6.63 2.33 5.70 2.38 4.01 READ-2016 4.61 0.03 3.85 4.88 4.34 0.73 4.34 1.09 3.80 0.80 4.22 0.63 Esposalles 0.61 0.05 5.86 19.99 1.91 2.23 5.95 4.91 2.16 3.37 2.13 2.86 ScribbleLens 4.72 0.06 4.38 16.24 4.81 2.17 5.06 4.04 4.16 2.71 4.31 2.24 CASIA-2 2.91 0.05 1.10 2.81 3.17 0.57 2.81 0.69 1.32 0.52 2.01 0.43 Eparchos 7.79 0.04 7.27 9.14 7.52 0.97 8.18 2.11 8.14 1.52 7.36 1.21 RIMES-2009 6.73 5.87 6.56 0.98 6.26 1.44 4.94 1.09 4.64 0.83 RIMES-2024 3.89 5.89 3.57 0.94 3.48 1.38 2.97 1.06 3.26 0.80 MAURDOR-C3 10.38 5.34 9.76 0.78 8.05 1.27 13.71 1.16 7.93 0.69 MAURDOR-C4 8.28 6.70 10.46 1.32 6.93 1.61 6.62 1.20 6.59 0.92 Average 5.31 9.54 5.33 1.17 5.36 2.36 4.97 1.81 4.51 1.37 # parameters 17.7 M 23.1 M 23.1 M 23.1 M 23.4 M 23.4 M
+We now provide a large evaluation at page level over all datasets and prediction approaches (ours and concurrent ones).
+We used the same architecture and training strategy for all the evaluated settings, i.e., we also train the concurrent approaches with the same framework for fair comparisons.
 We also provide results at line level, which corresponds to the pre-training stage (same encoder with CTC decoding) for comparison.
-However, let recall that line-level metrics only correspond to recognition on humanly-segmented images: it does not include the required segmentation and ordering stages that would lead to higher CER and prediction times.
-Results shown in Table VI-E highlight that it is generally beneficial to process documents at page level rather than line level.
+However, let us recall that line-level metrics only correspond to recognition on humanly-segmented images: it does not include the required segmentation and ordering stages that would lead to higher CER and prediction times.
+Results shown in Table 6.5 highlight that it is generally beneficial to process documents at page level rather than line level.
 A counter-example is the Esposalles dataset, where line recognition performs very well, demonstrating a certain ease of visual reading (single writer).
 In this particular case, it seems that the benefits of implicit language modeling provided by the transformer decoder are not sufficient to compensate for the added difficulty of learning the reading order.
-Another point is about the prediction time: the three proposed approaches are on par with the prediction speed of the Faster DAN.
-However, one can note that the Faster DAN speed improvement is highly dependent to the number of lines to be predicted in the document, contrary to the proposed approaches.
-Indeed, the Faster DAN is 16.3 times faster than the DAN for the BRESSAY dataset but 4.9 times faster for the CASIA-2 dataset, given that there are on average 30 lines per document for BRESSAY and 10 lines per document for CASIA-2.
-Regarding the CER performance now, the MT-DAN approach reaches new state-of-the art performance over 5 datasets.
-Also, it seems that the enlarged context provided by the Meta-DAN enables to stabilize the attention process, reducing the phenomenon of predictions loops (when the same text is predicted multiple times) for heterogeneous layouts as in the MAURDOC-C3 and MAURDOR-C4 datasets.
-It results in new state-of-the-art recognition performance when averaging on all datasets with an improvement of 0.46 point compared to MT-DAN and 0.80 point compared to DAN.
-WER results are given in Appendix C for comparison purposes.
+Focusing on the prediction time, all proposed approaches are on par with the prediction speed of the Faster DAN.
+However, one can note that the Faster DAN speed improvement is highly dependent on the number of lines to be predicted in the document, contrary to the proposed approaches.
+E.g., the Faster DAN is 16.3 times faster than the DAN for the BRESSAY dataset ($\approx 30$ lines/doc) but 4.9 times faster for the CASIA-2 dataset ($\approx$ 10 lines/doc).
+Regarding the CER, the MT-DAN reaches state-of-the-art performance on 4 datasets.
+Also, it seems that the enlarged context provided by the Meta-DAN stabilizes the attention process, reducing the looping phenomenon (when the same text is predicted multiple times) for heterogeneous layouts as in the MAURDOR datasets.
+It results in new state-of-the-art recognition performance when averaging on all datasets with an improvement of 0.46 points compared to MT-DAN and 0.80 points compared to DAN.
+Associated WER results are given in C for completeness.
+6.6 Comparison with the state of the art
 Comparison with state-of-the-art approaches (no external data, no external language model).
@@
 State-of-the-art line-level results are given for comparison.
-[b] Dataset Approaches of the literature CER (%) for our approaches (page level) Architecture Level CER (%) MT-DAN Meta-DAN BRESSAY [42] CRNN + attention line 3.62 2.33 2.38 [42] YOLO-v5 + Calamari line 6.01 [42] DANIEL page 3.55 CASIA-2 [48] CNN + attention line 2.34 1.32 2.01 [49] DenseNet + attention line 1.78 Eparchos [44] CNN + GRU line 4.53 8.14 7.36 Esposalles [21] DAN + NER line 0.48 2.16 2.13 [21] DAN + NER page 3.61 IAM [50] transformer line 3.99 4.51 4.79 [13] DAN page* 4.54 [17] DANIEL page* 4.38 MAURDOR-C3 [13] DAN page 8.62 13.71 7.93 [14] Faster DAN page 8.93 MAURDOR-C4 [13] DAN page 8.02 6.62 6.59 [14] Faster DAN page 9.88 READ-2016 [12] VAN line 4.10 3.80 4.22 [13] DAN paragraph 3.22 [13] DAN page 3.43 [14] Faster DAN page 3.95 [17] DANIEL page 4.03 RIMES-2009 [13] DAN page 4.54 4.94 4.64 [14] Faster DAN page 6.38 [17] DANIEL page 5.80 RIMES-2024 2.97 3.26 ScribbleLens [51] CNN + BLSTM line 6.47 4.16 4.31 * For IAM, the page level refers in the literature to pages for which the header including duplicated printed text was removed.
-VI-F Comparison with the state of the art
-We further compare the proposed approaches to the state of the art in Table VI-E.
-For fair comparison, we only compare to works following the same constraints, i.e., without external language model and that do not perform training on additional external dataset.
+[b] Dataset Approaches of the literature CER (%) for our page-level approaches Architecture Level CER (%) MT-DAN Meta-DAN BRESSAY [17] CRNN + attention line 3.62 2.33 2.38 [17] YOLO-v5 + Calamari line 6.01 [17] DANIEL page 3.55 CASIA-2 [47] CNN + attention line 2.34 1.32 2.01 [44] DenseNet + attention line 1.78 Eparchos [42] CNN + GRU line 4.53 8.14 7.36 Esposalles [40] DAN + NER line 0.48 2.16 2.13 [40] DAN + NER page 3.61 IAM [18] transformer line 3.99 4.51 4.79 [14] DAN page 4.54 [11] DANIEL page 4.38 MAURDOR-C3 [14] DAN page 8.62 13.71 7.93 [16] Faster DAN page 8.93 MAURDOR-C4 [14] DAN page 8.02 6.62 6.59 [16] Faster DAN page 9.88 READ-2016 [15] VAN line 4.10 3.80 4.22 [14] DAN page 3.43 [16] Faster DAN page 3.95 [11] DANIEL page 4.03 RIMES-2009 [14] DAN page 4.54 4.94 4.64 [16] Faster DAN page 6.38 [11] DANIEL page 5.80 RIMES-2024 2.97 3.26 ScribbleLens [20] CNN + BLSTM line 6.47 4.16 4.31
+We further compare the proposed approaches to the state of the art in Table 6.6.
+For fair comparison, we only compare to works following the same constraints, i.e., without an external language model and that do not perform training on additional external datasets.
 Among the 7 datasets for which there is a page-level evaluation in the literature, we reach new state-of-the-art CER on 4 of these datasets with the MT-DAN or the Meta-DAN.
 The proposed approaches also outperform state-of-the-art line-level recognition models on ScribbleLens and CASIA-2 datasets.
-On can note an exception with the Eparchos dataset, which follow a similar behavior compared to the Esposalles dataset, i.e., text recognition is better at line level than at page level.
-This can be explained by the very few available training data that prevent from fully benefiting from the full potential of attention-based recognition systems.
-Lastly, we provide a first evaluation for the newly released RIMES-2024 dataset with a CER of 2.97% reached by the MT-DAN.
-VI-G Multilingual recognition
-CER (in percentages) for different encoders on the test sets for line-level multilingual recognition.
-FCN-M ConvNeXtV2-N ConvNeXtV2-B Pre-training IN-21k IN-21k IN-21k Training Monolingual Multilingual Monolingual Multilingual Multilingual Multilingual IAM 4.84 6.40 5.54 9.57 6.34 5.31 BRESSAY 3.45 5.27 3.61 7.85 4.51 3.23 READ-2016 4.61 6.92 5.23 8.28 6.25 4.74 Esposalles 0.61 1.83 1.02 2.11 1.33 0.82 ScribbleLens 4.72 6.49 5.27 8.61 6.07 5.07 Average CER (%) 3.65 5.38 4.13 7.28 4.90 3.83 # parameters 17.6 M 16.0 M 88.8 M Average time (s) 0.03 0.01 0.03
-In this last experiment, we propose training the Meta-DAN in a multilingual setting, i.e., the model is trained on a set of datasets at once, and evaluated on the same datasets.
-We only focus on latin languages to see if learning one language can be helpful for learning another one when sharing part of the same alphabet.
-We compare the FCN-M encoder we used so far to two variants of the ConvNeXt-V2 architecture [52], which is a state-of-the-art CNN architecture for computer vision, pre-trained through self-supervised learning (masked auto-encoder) on ImageNet-21k (IN-21k).
-Regarding these ConvNeXt-V2 encoders, we made the following modifications for fair comparison and CTC requirements for pre-training (minimum input length with respect to the ground truth): 1) the last projection is set to output 256 feature maps to be compatible with the DAN decoder; 2) The ConvNeXt encoder includes five downsampling stages, each one dividing by two the horizontal and vertical axes.
-We removed two of them for the horizontal axis, leading to a final output shape for the image features of ($H_{f}=\frac{H}{32},W_{f}=\frac{W}{8},C_{f}=256$), which corresponds to the FCN-M encoder output shape.
-We first concentrate on the line-level pre-training stage in Table VI-G.
-As one can note, the FCN-M encoder is better suited for monolingual setting compared to the same-scale ConvNeXt-V2-N model, but we can observe the contrary for the multilingual setting.
-The results highlight that is mainly due to the self-supervised pre-training strategy on a large-scale IN-21k dataset: the overall CER is 2.38 point higher when training from scratch.
-Another interesting point is that, even with a larger model (ConvNeXt-V2-B, 4 times more parameters than FCN-M), the obtained results are still a bit less good as the monolingual approach with FCN-M encoder (3.83% of CER compared to 3.65%).
-We now compare these different encoders at page level with the Meta-DAN approach.
-To make training more efficient, we add an extra training step at line-level, leading to a 4-step training: 1) pre-training the encoder with CTC supervision through the line-level approach; 2) transfer learning to MT-DAN on line-level images; 3) further training the MT-DAN at page level with curriculum learning 4) transfer learning to Meta-DAN at page level without curriculum strategy.
-Table VI-G describes the obtained results, highlighting the need for an upscaled encoder to get close to monolingual results.
-Indeed, the ConvNeXt-V2-B encoder leads to an average CER of 4.62% on a multilingual setting, which is very close to the 4.45% obtained by the FCN-M with monolingual training.
-In addition, it only results in a small prediction time increase, from 1.57 second per page (FCN-M, monolingual) to 1.74 second per page (CN2-B, multilingual).
-Comparison of CER results (in percentages) at page-level with FCN-M and ConvNeXt-V2 (CN2) encoders for multilingual training using the Meta-DAN ($w=5$, $m=5$).
-Dataset Monolingual Multilingual FCN-M FCN-M CN2-N CN2-B BRESSAY 2.38 4.82 4.14 4.07 READ-2016 4.22 5.43 6.80 4.91 Esposalles 2.13 2.58 2.69 2.47 ScribbleLens 4.31 9.79 7.51 5.87 IAM 4.79 5.10 4.89 3.97 RIMES-2024 3.26 3.22 3.31 2.93 MAURDOR-C3 7.93 7.41 6.91 7.84 MAURDOR-C4 6.59 17.94 7.06 4.92 Average CER 4.45 7.04 5.41 4.62 Average time (s) 1.57 1.36 1.01 1.74 # parameters 23.2 M 23.2 M 21.5 M 94.3 M
-These results also demonstrates that the multilingual setting is only beneficial for similar languages regarding the evaluated models.
-Comparing monolingual FCN-M and multilingual CN2-B, one can notice that CER of all French/English datasets (IAM, RIMES-2024 and MAURDOR) are improved, while all datasets of historical documents (Esposalles, READ-2016 and ScribbleLens), which involves a specific period and language, are degraded.
-VII Discussion
-Through this extensive study on prediction strategies, we showed the robustness of the proposed Meta-DAN approach through many evaluation configurations (various datasets, multilingual recognition), reaching on-average state-of-the-art results.
-However, one can note that there is no one strategy that is better than the others on all datasets.
-We noticed that, no matter the prediction variant (Faster DAN or ours), it seems to have a beneficial impact on the loop effects that can occur during decoding, noticeable through the CER results of the Esposalles dataset (Table VI-E).
-We assume that for the Faster DAN strategy, this is due to the parallelization of the line recognition.
-Indeed, the loops mainly occur when switching from one line to another.
-For the approaches we proposed, this must be the result of the increased context modeling and the forced anticipation of the model to predict the following characters, which can be found on two consecutive lines.
+One can note an exception with the Eparchos dataset, which follows a similar behavior compared to the Esposalles dataset, i.e., text recognition is better at line level than at page level.
+This can be explained by the very few available training data that prevents from fully benefiting from the potential of attention-based models.
+6.7 Error analysis
+In this section, we propose a qualitative analysis of the errors made by the Meta-DAN.
+To this end, some examples of predictions and associated ground truths are shown in Figure 6.7 to illustrate the different kinds of errors that can occur.
+The most common error is linked to visual ambiguities.
+For instance, the Meta-DAN predicted an "m" instead of "rr" for the word "arrived" (Figure 6.7) or and "n" instead of an "r" in the word "garida" (Figure 6.7).
+This shows that, although the proposed contributions improve context modeling, it remains limited: the model sometimes tends to rely more on visual characteristics than on linguistic context.
+We assume that this is due to the lack of training data in the datasets that prevents the model from really "understand" the text it predicts, leading to the prediction of "dest" instead of "de St" (Figure 6.7), due to the spacing between those two words that is not clearly visible in the image.
+In other words, it suffers from a lack of linguistic knowledge, but it also lacks knowledge of the world around us to clear up certain ambiguities.
+Another kind of error comes from the attention-based prediction process itself.
+In this way, the same character might be predicted twice (Figure 6.7), or some characters are missed, as in Figure 6.7 ("P.S"), generally due to specific layouts not seen during training.
+In rare cases, some characters may have been predicted in the wrong order and partially missed, as "Davenun" predicted instead of "Dans un avenir" (Figure 6.7).
+As mentioned previously, another failure case due to the attention mechanism is related to loops, when the model predicts the same part of text, generally a whole text line, multiple times.
+Although rare, predictions may be considered false, wrongly, when the ground truth contains errors as in Figure 6.7: the Meta-DAN predicted "d’en", as visible in the input image, but the ground truth is "de".
+CASIA-2
+IAM
+Esposalles
+MAURDOR-C3
+RIMES-2009
+Visualization of errors made by Meta-DAN on test samples of various datasets.
+For visibility, page images were cropped and associated with Ground Truth (GT) and Prediction (P).
+Errors are underlined in red.
+7 Discussion
+Through this extensive study on prediction strategies, we showed the robustness of the proposed Meta-DAN on many HTR dataset configurations, reaching on-average state-of-the-art results.
+However, one can note that there is no one strategy that is better on all datasets.
+We noticed that, no matter the prediction variant, it seems to have a beneficial impact on the loop effects, noticeable through the CER results of the Esposalles dataset (Table 6.5).
+We assume that for the Faster DAN, this is due to the parallelization of the line recognition, since loops mainly occur when switching from one line to another.
+For the proposed approaches, this must be the result of the increased context modeling and the forced anticipation of the model to predict the following characters, which can be found on two consecutive lines.
 But contrary to the Faster DAN, the Meta-DAN does not suffer from "catastrophic" errors.
-Indeed, any missed or duplicated character in the first decoding stage of the Faster DAN lead to discard or duplicate a whole text line.
-While rare, this can severely impact the recognition performance, also explaining the recognition gap.
-Although we limit ourselves to compare the approaches for character-level predictions to reach state-of-the-art results with few training data, it has to be noted that all approaches could also be used in combination with BPE or N-gram tokenization to further reduce the prediction time.
-VIII Conclusion
-In this paper, we proposed and evaluated two new prediction strategies for end-to-end full-page text recognition: windowed queries and multi-token predictions.
+Indeed, any missed or duplicated character in the first decoding stage of the Faster DAN leads to discarding or duplicating a whole text line.
+While rare, this can severely impact the recognition performance.
+Regarding the prediction time, it has to be noted that the Meta-DAN, while competitive, does not outperform the Faster DAN in this respect, on average.
+Indeed, although the hyperparameters $m$ and $w$ can be arbitrarily chosen, the experiments showed that in practice, they must be set to a rather low value to preserve competitive recognition performance.
+Thus, the current state-of-the-art prediction speed for end-to-end page-level text recognition remains an issue for real-time applications.
+Although we limit ourselves to comparing the approaches for character-level predictions to reach state-of-the-art results with few training data, all approaches could be combined with BPE tokenization to further reduce the prediction time.
+8 Conclusion
+In this paper, we proposed two new prediction strategies for end-to-end full-page text recognition: windowed queries and multi-token predictions.
 We show that the Meta Document Attention Network approach, combining both strategies, reaches new on-average state-of-the-art results on many handwritten datasets while reaching competitive prediction times.
-In addition, this approach enables flexible balance between prediction time and recognition performance by adapting the associated hyperparameters to adapt to the use case.
-While we only evaluated the Meta-DAN for HTR, the proposed approach is generic enough to be adapted to any autoregressive transformer-based system.
-Lastly, we highlight that using a generic self-supervised pre-trained encoder is an effective way to bridge the performance gap between monolingual and multilingual settings.
+This approach offers a flexible balance between prediction time and recognition performance by adapting the associated hyperparameters to the given use case.
+While we only evaluated the Meta-DAN for HTR, the proposed approach could be applied to any autoregressive transformer-based system.
+However, there is still room for improvement, and some issues remain to be solved regarding end-to-end full-page text recognition.
+The evaluation protocol of state-of-the-art approaches is based on training a specific model on a single dataset at a time.
+How to scale them to handle multilingual content, multiple reading orders (mixing Latin and Arabic languages, for instance), and highly heterogeneous layouts remains to be explored.
+These limited training configurations prevent the models from acquiring large, generic linguistic knowledge, leading to some failure cases when the visual content is degraded or ambiguous.
+Another point is about the prediction time, which remains an issue for real-time applications, as it is still in the order of magnitude of one second per document.
+Finally, current state-of-the-art approaches consider a single reading order per document, but this is not realistic for very complex layouts (e.g., schemas, maps), for which several reading orders could be considered as humanly correct.
+Future works should focus on adapting metrics and losses to allow multiple reading orders for those cases.
 Acknowledgment
 This work was granted access to the HPC resources of IDRIS under the allocation 2024-AD011012155R2 made by GENCI.
-Experiments presented in this paper were also partly carried out using the Grid’5000 testbed, supported by a scientific interest group hosted by Inria and including CNRS, RENATER and several Universities as well as other organizations (see https://www.grid5000.fr).
-APPENDIX
-VIII-A Scaling up the DAN encoder
-We noticed that the DAN lacked of some feature extraction capability when evaluated on challenging datasets such as BRESSAY which presents some crossed-out text with annotations.
-We opt for scaling up the DAN encoder only in order to increase the performance while preventing a significant prediction time increase.
-Indeed, the encoding part of the model is only used once at inference (it is excluded from the iterative decoding stage); so the computation involved are not such significant compared to decoding.
-The proposed encoder is referred to as FCN-M, and the original one as FCN-S.
-The proposed modification mainly consists in doubling the number of feature maps per block while preserving the last one which condition the size of the decoder, as detailed in Table VIII-A.
-We use the notation of the original paper [12] where ConvBlock is a Convolution Block made up of three convolutions, instance normalization and dropout, and DSCBlock is equivalent except for the convolutions that are replaced by Depthwise Separable Convolutions [36].
-Composition of the two versions of the encoder.
-ConvBlock(X) and DSCBlock(X) indicate the number of feature maps X returned by each convolutional layer.
-FCN-S [12] FCN-M (ours) Output size layers ConvBlock(16) ConvBlock(32) $H\times W$ ConvBlock(32) ConvBlock(64) $\frac{H}{2}\times\frac{W}{2}$ ConvBlock(64) ConvBlock(128) $\frac{H}{4}\times\frac{W}{4}$ ConvBlock(128) ConvBlock(256) $\frac{H}{8}\times\frac{W}{8}$ ConvBlock(128) ConvBlock(512) $\frac{H}{16}\times\frac{W}{8}$ ConvBlock(128) ConvBlock(512) $\frac{H}{32}\times\frac{W}{8}$ DSCBlock(128) DSCBlock(512) $\frac{H}{32}\times\frac{W}{8}$ DSCBlock(128) DSCBlock(512) $\frac{H}{32}\times\frac{W}{8}$ DSCBlock(128) DSCBlock(512) $\frac{H}{32}\times\frac{W}{8}$ DSCBlock(256) DSCBlock(256) $\frac{H}{32}\times\frac{W}{8}$ params 1.7 M 17.6 M
-We evaluate these modifications on READ-2016, CASIA-2, BRESSAY and IAM datasets at line level, and at page level with the DAN.
-Results are shown in Table VIII-A.
-As one can note, the modified encoder improves the recognition results (or is equivalent) on all the evaluated datasets, whether it is at line or page level.
-The prediction time is only slightly impacted, we assume this is mainly due to the complexity of learning the reading order.
-For CASIA-2 and IAM, the prediction time is lower, which can be explained by a reducing of the loop effect (predicting several times the same text part).
-But for READ-2016 and BRESSAY, the prediction time is higher, which can be the result of the additional computations due to the bigger encoder, and to a better recognition (missed text parts are now well recognized, adding extra decoding iterations).
-The most negative impact is on the READ-2016 dataset with a 6% increase of the prediction time.
-These results also enhance what have been found in [13]: the page-level DAN outperforms its line-level counterpart while they are based on the same encoder.
-Impact of the encoder changes at line (pre-training stage) and page levels.
-CER are expressed in percentages, and times in seconds.
-Dataset Encoder FCN (line) DAN (page) CER Time CER Time READ-2016 FCN-S 4.82 0.03 3.82 4.61 FCN-M 4.61 0.03 3.85 4.88 CASIA-2 FCN-S 5.33 0.05 1.35 2.94 FCN-M 2.91 0.05 1.10 2.81 Bressay FCN-S 6.11 0.02 4.53 24.28 FCN-M 3.45 0.02 2.43 24.65 IAM FCN-S 5.76 0.03 4.53 3.57 FCN-M 4.84 0.03 4.23 3.39
-VIII-B Datasets
-Line-level splits
-Table VIII-B details the split for training, validation and test used for the different datasets regarding the experiments at line level.
-Datasets splits at line level.
-Dataset Train Validation Test IAM 6,482 976 2,915 READ-2016 8,367 1,043 1,140 Esposalles 2,328 742 757 ScribbleLens 4,408 375 563 CASIA-2 36,305 5,465 10,449 Eparchos 1,512 380 380 BRESSAY 19,561 4,608 5,915
+Experiments were also carried out using the Grid’5000 testbed, supported by a scientific interest group hosted by Inria and including CNRS, RENATER and several Universities as well as other organizations (see https://www.grid5000.fr).
+Appendix A Datasets
 Example images
-Figure VIII-B provides document image examples from the 10 evaluated datasets. As one can note, these datasets cover many page layouts, visual aspects, writing styles, periods and languages.
+Figure A provides document image examples from the 10 evaluated datasets. As one can note, these datasets cover many page layouts, visual aspects, writing styles, periods, and languages.
 MAURDOR-C3
@@
 Dataset samples from test sets. The MAURDOR-C3 and MAURDOR-C4 datasets provide widely heterogeneous documents in terms of layout, while the other datasets are quite homogeneous.
-VIII-C WER results for the comparison of prediction strategies.
-Table VIII-C provides the WER scores associated to experiments related to Section VI-E, i.e., when relying on monolingual evaluation and the FCN-M encoder.
-Interestingly, the WER results do not follow exactly the same trend as that of the CER, by exhibiting a slight advantage for the Faster DAN with reaches an on-average WER of 13.73%, compared to 14.22% for the Meta-DAN.
-There is no clear explanation for this observation.
-However, it still remains lower than the DAN WER which reaches 15.15%.
+Line-level splits
+Table A details the splits for training, validation, and test used for the different datasets regarding the experiments at line level.
+Datasets splits at line level.
+Dataset Train Validation Test IAM 6,482 976 2,915 READ-2016 8,367 1,043 1,140 Esposalles 2,328 742 757 ScribbleLens 4,408 375 563 CASIA-2 36,305 5,465 10,449 Eparchos 1,512 380 380 BRESSAY 19,561 4,608 5,915
+Appendix B Scaling up the DAN encoder
+We noticed that the DAN lacked some feature extraction capability when evaluated on challenging datasets such as BRESSAY, which presents some crossed-out text with annotations.
+We opt for scaling up the DAN encoder only in order to increase the performance while preventing a significant prediction time increase.
+Indeed, the encoding part of the model is only used once at inference (it is excluded from the iterative decoding stage); so the computations involved are not such significant compared to decoding.
+The proposed encoder is referred to as FCN-M, and the original one as FCN-S.
+Composition of the two versions of the encoder.
+ConvBlock(X) and DSCBlock(X) indicate the number of feature maps X returned by each convolutional layer.
+FCN-S [15] FCN-M (ours) Output size layers ConvBlock(16) ConvBlock(32) $H\times W$ ConvBlock(32) ConvBlock(64) $\frac{H}{2}\times\frac{W}{2}$ ConvBlock(64) ConvBlock(128) $\frac{H}{4}\times\frac{W}{4}$ ConvBlock(128) ConvBlock(256) $\frac{H}{8}\times\frac{W}{8}$ ConvBlock(128) ConvBlock(512) $\frac{H}{16}\times\frac{W}{8}$ ConvBlock(128) ConvBlock(512) $\frac{H}{32}\times\frac{W}{8}$ DSCBlock(128) DSCBlock(512) $\frac{H}{32}\times\frac{W}{8}$ DSCBlock(128) DSCBlock(512) $\frac{H}{32}\times\frac{W}{8}$ DSCBlock(128) DSCBlock(512) $\frac{H}{32}\times\frac{W}{8}$ DSCBlock(256) DSCBlock(256) $\frac{H}{32}\times\frac{W}{8}$ # params 1.7 M 17.6 M
+The proposed modification mainly consists in doubling the number of feature maps per block while preserving the last one, which conditions the size of the decoder, as detailed in Table B.
+We use the notation of the original paper [15], where ConvBlock is a Convolution Block made up of three convolutions, instance normalization, and dropout.
+DSCBlock is equivalent except for the convolutions that are replaced by Depthwise Separable Convolutions [9].
+Impact of the encoder changes at line (pre-training stage) and page levels.
+CER are expressed in percentages, and times in seconds.
+Dataset Encoder FCN (line) DAN (page) CER Time CER Time READ-2016 FCN-S 4.82 0.03 3.82 4.61 FCN-M 4.61 0.03 3.85 4.88 CASIA-2 FCN-S 5.33 0.05 1.35 2.94 FCN-M 2.91 0.05 1.10 2.81 Bressay FCN-S 6.11 0.02 4.53 24.28 FCN-M 3.45 0.02 2.43 24.65 IAM FCN-S 5.76 0.03 4.53 3.57 FCN-M 4.84 0.03 4.23 3.39
+We evaluate these modifications on READ-2016, CASIA-2, BRESSAY, and IAM datasets at line level, and at page level with the DAN.
+Results are shown in Table B.
+As one can note, the modified encoder improves the recognition results (or is equivalent) on all the evaluated datasets, whether it is at line or page level.
+The prediction time is only slightly impacted; we assume this is mainly due to the complexity of learning the reading order.
+For CASIA-2 and IAM, the prediction time is lower, which can be explained by a reduction of the loop effect (predicting several times the same text part).
+But for READ-2016 and BRESSAY, the prediction time is higher, which can be the result of the additional computations due to the bigger encoder, and to better recognition (missed text parts are now well recognized, adding extra decoding iterations).
+The most negative impact is on the READ-2016 dataset with a 6% increase in the prediction time.
+These results also enhance what has been found in [14]: the page-level DAN outperforms its line-level counterpart while they are based on the same encoder.
+Appendix C WER results comparing the prediction strategies.
+Table C provides the WER scores associated with the experiments related to Section VI-E, i.e., when relying on the FCN-M encoder.
 WER results (in percentages) on the test sets with FCN-M encoder.
@@
 Dataset Line-level Page-level FCN DAN Faster DAN W-DAN MT-DAN Meta-DAN IAM 15.83 14.00 13.42 16.28 14.96 15.78 BRESSAY 12.12 7.85 8.03 9.72 8.28 8.46 READ-2016 20.41 14.69 14.83 16.90 14.90 16.03 Esposalles 2.61 6.60 2.73 6.63 3.09 3.09 ScribbleLens 20.38 19.66 18.64 21.19 18.92 19.71 Eparchos 26.29 22.36 22.36 24.71 23.39 23.69 RIMES-2009 14.35 12.99 15.12 12.67 12.96 RIMES-2024 8.72 8.30 9.35 8.00 8.72 MAURDOR-C3 28.43 19.82 20.71 27.46 20.23 MAURDOR-C4 14.80 16.19 13.50 13.15 13.48 Average 15.15 13.73 15.41 14.48 14.22
+Interestingly, the WER results do not follow exactly the same trend as that of the CER, by exhibiting a slight advantage for the Faster DAN, which reaches an on-average WER of 13.73%, compared to 14.22% for the Meta-DAN.
+There is no clear explanation for this observation.
+However, it still remains lower than the DAN WER, which reaches 15.15%.
+This would suggest that the Faster-DAN is more reliable when its initial predictions are correct, leading to perfectly recognized words (lower WER), but that bad starts cause cascading errors, leading to higher CER, while the proposed Meta-DAN is more constant in character recognition performance.